### v1

# Rethinking Programmed I/O for Fast Devices, Cheap Cores, and Coherent Interconnects

Abstract.
Conventional wisdom holds that an efficient interface between an OS running on a CPU and a high-bandwidth I/O device should be based on Direct Memory Access (DMA), descriptor rings, and interrupts: DMA offloads transfers from the CPU, descriptor rings provide buffering and queuing, and interrupts facilitate asynchronous interaction between cores and device with a lightweight notification mechanism.
In this paper we question this wisdom in the light of modern hardware and workloads, particularly in cloud servers. We argue that the assumptions that led to this model are obsolete, and in many use-cases use of programmed I/O, where the CPU explicitly transfers data and control information to and from a device via loads and stores, actually results in a more efficient system.
We quantitatively demonstrate these advantages using three use-cases: fine-grained RPC-style invocation of functions on an accelerator, offloading of operators in a streaming dataflow engine, and a network interface targeting for serverless functions. Moreover, we show that while these advantages are significant over a modern PCI Express (PCIe) peripheral bus, a truly cache-coherent interconnect offers significant additional efficiency gains.
1. Introduction
There is a current consensus on how CPUs should communicate with modern, high-performance devices like network interface adaptors (NICs), storage controllers, Graphics Processing Units (GPUs), and other computational accelerators: for the most part, this is achieved using DMA and rings of descriptors to decouple CPU and device as much as possible. On modern PCIe interconnects, this consensus results in impressive bandwidth for applications which transfer data in large chunks.
However, in this paper we question this consensus in the light of several factors: (1) important workloads involving irregular access and fine-grained interaction between CPU and device, (2) increasing core counts in servers, (3) data center applications where many machines are executing the same application on most of their cores, and (4) the emergence of cache-coherent peripheral interconnects, particularly those like CXL.cache 3.0 (CXL Consortium, 2024a) in which devices are first-class participants in the coherence protocol.
We argue that these factors make Programmed I/O (PIO), where CPU software performs I/O by directly reading and writing device registers, a better mechanism for data transfer in some important scenarios than one based on DMA.
We use a real hardware platform which implements a coherent interconnect between a server-class CPU and a large FPGA to investigate the trade-offs between DMA using descriptor rings, PIO directly over a PCIe interconnect, and PIO over a cache-coherence protocol.
We show that, for small (around 1KiB or less) transfers, PIO posted writes over PCIe significantly outperform DMA, although reads are less efficient since, unlike writes, they cannot be pipelined and so incur PCIe’s significant round-trip latency penalty. However, a full cache coherence protocol which avoids PCIe significantly outperforms both DMA and PIO over PCIe.
We go further, however, and show how a conventional MESI-like coherence protocol can be exploited in novel ways by intelligent devices which can interact with an unmodified CPU at the coherence message level. By relaxing traditional coherence protocol assumptions (for example, the independence of cache lines), we can achieve dramatically more efficient communication between a CPU and a device.
In the following section, we provide background and motivation for this work, including the factors that led to the descriptor-based DMA model of I/O, and why it is time to question the underlying assumptions in that model. In Section 3 we describe and motivate the experimental hardware platform we use for our evaluation, and also provide calibrating performance comparisons with PC data center servers.
In Section 4, we discuss why true cache-coherent device interconnects are fundamentally different from today’s PCIe platforms, and present a set of efficient protocols for passing messages with low latency between software running on an unmodified CPU, and a smart device which has low-level access to the coherence protocol. In Section 4.5, we go on to present the implementation of some of these protocols on our experimental platform.
Section 5 then compares a traditional DMA-based approach with PIO over PCIe and our new coherence-based protocols using three different use-cases: (1) lightweight, synchronous local invocation of functionality on a computational accelerator, (2) I/O to and from an intelligent network interface, and (3) hardware offload of a distributed dataflow stream processor to an FPGA. We survey related work in Section 6 and conclude with Section 7.
2. Background and Motivation
Our work in this paper is motivated by the interaction between modern trends in platform interconnect and also the changing ways in which these interconnects are used, particularly in data center and cloud servers.
2.1. Background: hardware/software interaction
In early machines, the CPU interacted with peripheral devices exclusively using programmed I/O (PIO): the device exposed hardware registers mapped into the processor’s address space. Both data transfer and control of the device involved reads and writes by the CPU to these registers.
This model made sense when devices were extremely simple, particularly when augmented by the use of device interrupts to obviate the need for the CPU to poll status registers on the device waiting for an I/O operation to complete. However, the model still requires a synchronous rendezvous between the CPU and device, and moreover requires all data to be transferred via CPU registers (since the CPU is copying data between device registers and main memory).
DMA removes this limitation by allowing the device itself or a dedicated DMA controller to copy the data while the CPU continues to execute other code, providing parallelism and partly decoupling the CPU and device. Further decoupling is provided by descriptor rings: producer/consumer buffers of I/O requests in memory represented as a queue of operations, and read and written by both the CPU and the DMA-capable device. In this case, interrupts need only be used when the queue of descriptors becomes full or empty.
Almost every high-speed device today uses essentially this technique, with minor variations such as how the descriptor queues are formatted, where they are stored, and whether the head and tail of the queue is identified by registers or flags in the descriptors themselves.
This model is optimized for throughput – it is most efficient when a large volume of data must be transferred.
2.2. Trends in interconnects and devices
[…]
Recent trends in accelerator design, however, have led to proposed standards for cache coherent peripheral interconnects, and the related idea of extending processor cache coherence to heterogeneous devices: accelerators, networking interfaces, and storage controllers (Stuecheli et al., 2018; CCIX Consortium and others, 2024; CXL Consortium, 2024b; Nvidia, 2024; AMBA® CHI Architecture Specification, 2024). The more far-reaching of these proposals are fully symmetric, i.e. devices are first-class participants in a distributed directory-based cache coherence protocol.
Communication in software between peer modes in a cache-coherent system is typically very different from using descriptor-based DMA. FastForward (Giacomoni et al., 2008a), Barrelfish’ UMP (Baumann et al., 2009a), Concord (Iyer et al., 2023), and Shinjuku (Kaffes et al., 2019) adopt a much more direct approach to sending small messages with low latency, e.g. by exploiting the cache coherence protocol to transfer lines between caches on demand, and local polling to provide synchronization. Recently, cache coherence between NUMA CPUs has been used to simulate communicating between software and a hypothetical cache-coherent NIC (Schuh et al., 2024).
2.3. “New” workloads
Without denying the importance of throughput-oriented data transfer between devices and cores for workloads like current machine learning training and inference, it is worth stepping back and questioning many of the assumptions that led us here, and what other workloads might be a poor fit today for this model.
Any workload involving fine-grained, frequent interaction between the CPU and another device incurs a significant latency penalty when using DMA for data transfer. Many irregular workloads have this property (Gioiosa et al., 2017; Bean, 2016; Kashimata et al., 2019; Zeng and Prasanna, 2020).
Moreover, data center messaging (essentially, network Remote Procedure Calls (RPCs)) exchange many small messages (Karandikar et al., 2021; Lazarev et al., 2021). While end-to-end latency is mostly network propagation time in this case, latency incurred in the end system is a good proxy for resources (such as CPU cycles) wasted before invoking the application code, and in marshaling and sending any reply.
There are also several assumptions underlying the DMA-based communication model which are worth revisiting. In particular, CPU execution and I/O were decoupled because CPU cycles were precious (there being few cores), and the CPU had plenty of different tasks to work on at any time.
This is no longer true. Server CPUs have had many cores for some time, and moreover when running data center applications they are often dedicated to a single application, sometimes with miscellaneous functions moved to a small subset of the cores (Marty et al., 2019).
Another example: for throughput-oriented workloads DMA has evolved to efficiently transfer data to and from main memory without polluting the CPU cache. However, for small message applications, it is important that almost all the data gets into the right CPU cache as quickly as possible.
This leads us to rethink the use of the DMA model, and whether directly involving the CPU in data transfers to and from devices might be a better approach for many scenarios on modern hardware. Others have also suggested this in some cases (Larsen and Ben Lee, 2015; Chaudhari et al., 2017), but we also push the question much further: given a future cache-coherent interconnect, what new ways of using Programmed I/O might become compelling for data center applications.
3. Experimental platform
We investigate the modern trade-offs between PIO, DMA, and cache-coherent interconnects using the Enzian research computer (Cock et al., 2022). Enzian is useful as a placeholder for future, interoperable, cache-coherent interconnects like CXL.
We use a real hardware platform rather than a simulator. Simulation would be appropriate for detailed quantitative comparisons of known techniques, but in this paper we explore unconventional uses of cache coherence and so want to establish not only that our techniques are performant, but that they can practically be implemented in a real system.
An Enzian machine can be thought of as a two-socket NUMA server, where one socket houses a Marvel Cavium ThunderX-1 CPU, with a Xilinx XCVU9P FPGA in the other.
The CPU is a 48-core ARMv8 processor running at 2.0GHz, and 128GiB of 2133MT/s DDR4 memory spread across 4 memory controllers. Each core has a 32KiB, 32-way associative, write-through, physically indexed, physically tagged, L1 data cache. These are connected to a 16MiB, write-back, 16-way associative, shared L2 last-level cache. Hardware keeps the L1 caches coherent with the L2, and so in ARM architecture parlance the point of coherence is the L1 cache, whereas the point of unification is the L2 cache write-through. The cache-line size is 128 bytes – note that this is double the conventional 64-byte line size.
The CPU and FPGA are connected by the CPU’s native inter-socket cache coherent protocol, the Enzian Coherence Interface (ECI), which Enzian implements on the FPGA. It is a MOESI-like directory based protocol, and the physical address space is statically partitioned between the two NUMA nodes (CPU and FPGA). It uses 24 bidirectional lanes (organized into 2 12-lane links), each running at about 10Gb/s, for a theoretical inter-socket bandwidth of about 30GiB/s.
Both nodes have PCIe interfaces: the ThunderX-1 has a PCIe Gen.3 x8 interface, while the FPGA has a PCIe Gen.3 x16 interface. In this work we connect these two interfaces with a loopback cable; this essentially replicates the case of an PCIe accelerator card and allows us to compare PCIe with ECI in our experiments.
3.1. Characterizing the platform
Enzian provides a real hardware platform with server-class performance to evaluate different forms of device communication. However, direct comparisons with modern server platforms is not straightforward: the latter do not offer a full cache coherence protocol between the CPU and devices, but do provide more recent and therefore higher-performance cores, memory systems, and PCIe interconnects.
Therefore, in order to calibrate against a modern PC for both PIO and DMA over PCIe, we present two benchmark comparisons of Enzian with a modern PC (Intel Core i7-7700 3.6GHz Kaby Lake, PCIe Gen.3 x16) with cache line size of 64 bits connected to an AMD Xilinx Virtex UltraScale+ VCU118 card, using the same “write then read” experiment detailed in Section 5.1. The VCU118 uses the same FPGA as Enzian, albeit a slightly slower speed gauge.
3.1.1. DMA performance over PCIe
The first experiment uses DMA, using the Xilinx XDMA IP and associated descriptor-based protocol to transfer data between the CPU and the FPGA over PCIe. Figure 1 shows results for various data transfer sizes, running the Xilinx driver in both interrupt-driven and polled modes.
Figure 1. Comparison of XDMA latency for RPC workload between Enzian and PC
We observe that, using PCIe, exchanging messages using XDMA is about 3 times as fast on the PC server than on Enzian, while the difference between interrupt-driven and polling performance is much less significant. The latency of a single transaction is pretty constant on both platforms up to the PCIe transaction size limit of 4KiB, and then increases.
The difference in performance between the two architectures is due to several factors. For very small transfers, the CPU overhead of setting up descriptors dominates, and an x86 core is simply faster than the slower (and mostly in-order) ThunderX-1 core. In addition, the memory system is somewhat faster on the PC. The factor of 2 difference in PCIe bandwidth between the two platforms does not appear to be a factor in these experiments.
3.1.2. PIO performance over PCIe
Our second calibration experiment involves the CPU itself doing PIO reads and writes over PCIe to the FPGA.
We experimented with several different optimizations in this experiment, in particular the use of write combining and vector instructions. We present the best figures we could obtain on both platforms.
Figure 2. Comparison of PIO latency for RPC workload between Enzian and PC
Figure 2 shows the results. For transaction sizes above 32 bytes, the PC is about twice as fast as Enzian over the PCIe cable between CPU and FPGA. Here, the factor of 2 difference in PCIe bandwidth is responsible for the latency difference, in particular for reads (which are limited to 128 bits on both platforms). Writes, in contrast, are combined and highly pipelined at the PCIe interface.
4. Programmed I/O over coherent interconnects
In this section we explore what programmed I/O might look over coherent interconnects. It is of course possible for a CPU to simply copy data over a coherent interconnect as in conventional programmed I/O. This requires careful management of the processor’s cache hierarchy with explicit cache flush and invalidate instructions, but can be appropriate when the underlying hardware standard itself mandates a cache on the device, as with OpenCAPI 4.0 (Stuecheli et al., 2018; Allison et al., 2020).
That said, here we are more interested in how a coherent interconnect can best be used for low-latency, efficient movement of data between CPU registers and a device, including cases where the conventional notion of the cache as a view over memory contents is not appropriate (such as data arriving over a network).
We scope the question further as follows: we consider the case where the CPU cores, or their Last-Level Cache (LLC), communicate with devices via a MESI-like (Nagarajan et al., 2020), message-based coherence protocol with stable states corresponding to Modified (dirty), Shared (clean), Exclusive (clean), and Invalid, using a distributed cache directory. Note that this means that each address or cache line has a home node, which maintains the directory entry for the line.
Moreover, we assume a symmetric protocol: each device is a first-class participant in the protocol with a subset of physical addresses homed at it, and the device therefore maintains directory state for these lines.
While these assumptions are rare in hardware today, they are true both of the proposed CXL.cache standard (CXL Consortium, 2024a) and the Enzian platform we use for experiments.
4.1. Baseline: naively communicating with coherence
Our starting point is existing techniques which use cache coherence to perform inter-core communication, for example  (Giacomoni et al., 2008a) or Barrelfish UMP  (Baumann et al., 2009a). The basic idea is shown in Figure 3; to software, the sender writes data to a cache line on which the receiver is spinning (reading in a tight loop). Initially, the line is in (say) Shared state in both caches. When the sender writes the line (typically in one go due to the core’s write-combining buffer), it transitions briefly to Modified state, before being fetched by the receiver over the interconnect and reverting to Shared.
Figure 3. Simple inter-core communication over coherence
This can be highly efficient in software, particularly when deployed as a ring buffer of cache lines, and results in very little interconnect traffic (since spinning is usually on a local copy of the line). One can easily replace one core with a device in this scheme (with suitable state machine) to obtain an I/O protocol. The downside is, of course, the CPU cost of spinning, both in wasted cycles and the inability to switch to another task.
The design of protocols like FastForward is based on the idea that each endpoint can only manipulate its local cache using software instructions: loads, stores, etc. Moreover, it is assumed (reasonably) that flush or invalidate instructions should be avoided because they might be slow, and might require kernel privilege.
This is not the case when one of the endpoints is a device with message-level access to the cache coherence protocol, and possibly maintaining its own directory state.
4.2. The implications of low-level access
Unlike software running on a core behind a cache, a device participating in a coherence protocol has visibility over more state and events, and can also initiate more operations. Note that there is no need for a cache per se on the device at all, indeed this is usually inappropriate.
First, the device receives messages from the CPU cache that request lines in Shared or Exclusive state, or which request that the caching state on the be downgraded (e.g. from Exclusive to Shared, or from Shared to Invalid). It is therefore able to react internally to such events as it chooses, as long as it does not violate the protocol.
Second, the device can issue such requests itself at any time, allowing more fine-grained control of the state of a line in the CPU’s cache.
Third, for lines which are homed at the device, the device’ directory maintains explicit information about the cache line state at the CPU and other nodes as well as locally.
Fourth, unlike a conventional cache, the device does not have to respond immediately to every request from the CPU’s cache. Instead, it can choose to delay the response (blocking the requesting core) until some other operations complete. Care must be taken to respond before any hardware-imposed timeouts, but these are typically generous (hundreds of milliseconds on the ThunderX-1, for example) and can be worked around in software (e.g. by sending a “try again” response within the timeout).
Finally, the device can choose to interpret events like remote requests as more than simply loads or stores to the line concerned. This last insight is important because it, in combination with the previous observation, it allows us to construct efficient communication mechanisms which rely on a combination of the device’s richer view of the protocol and the software on the CPU adhering to additional conventions to communicate channel semantics to the device.
Put simply, requests to specific cache lines can be used to signal particular higher-level operations to the device, much as register reads and writes to traditional devices frequently have semantics very different from memory access.
This is an oversimplification, and we discuss the subtleties that arise below. First, however, we present four related communication protocols, layered above MESI-like coherence, which provide for efficient, blocking communication between a conventional CPU and a coherence-aware device.
(a) CPU-initiated, device-homed
(b) Device-initiated and homed
(c) CPU-initiated, CPU-homed
(d) Dev.-initiated, CPU-homed
Figure 4. Protocol variants for efficient CPU-device messaging
4.3. Communication protocol variants
We present 4 different but similar communication protocols (shown in Figure 4) which vary based on (a) the CPU or the device initiates the message transfer, and (b) the cache lines used are homed on the CPU or device. All provide a synchronous RPC: a transfer of a single cache-line-sized message, followed by a single cache-line response. All use 2 cache lines, which we designate A and B. These protocols can be generalized to larger payloads, and latency will grow linearly with the payload size. Interestingly, they can be made slightly more efficient by using more cache lines, but quickly become complex. We focus on the simple 2-line case in this paper.
The protocols use the same techniques and insights, and are best described using the example of a CPU-initiated RPC using addresses homed on the device (4(a)). An implementation of this protocol is used for the evaluation in Section 5.1.
  1. (1)
    The protocol starts with line B in Exclusive state in the CPU cache, and line A Invalid. To initiate an operation, the CPU writes the request data from its registers into line B, which transitions to Modified. This happens without communication with the device, since B was previously in Exclusive.
  2. (2)
    To signal to the device that line B holds a new request, the CPU then executes a load to line A. Since A is Invalid, the cache requests A in Shared from the device. This violates the typical convention that cache line operations are independent, but the device is free to interpret this request for A as a signal from the CPU that B holds a request.
  3. (3)
    The device does not respond immediately to the request for A, thereby stalling the load and blocking the requesting core. Instead, it requests B from the CPU’s cache, in Exclusive. The interconnect therefore transfers the request data to the device, and line B duly transitions to Invalid in the CPU’s cache.
  4. (4)
    The device reads the data it has received in B, and computes a result on it, which is to be returned to the CPU. It returns this using the response to the CPU’s request for cache line A.
  5. (5)
    The CPU cache receives line A’s new data and the CPU unblocks, immediately loading the first word of the RPC response into a register.
As described, the resulting state is that A is now Invalid and B is Shared in the CPU cache. However, in some deployed MESI-like protocols (including the Enzian interconnect we use), the device is free to return the data stipulating that it is not Shared but now Exclusive to the CPU. This final feature leaves the cache in the state it started in, with the roles of A and B reversed, ready for the next transaction.
The whole exchange takes 2 round trips on the coherent interconnect.
The other 3 protocols work similarly. For example, 4(b) shows the analogous case where the device initiates the request to the CPU, but still using lines homed on the device. This is used for the NIC experiment in Section 5.3.
The remaining two cases, where the cache lines are homed on the CPU, are more complex to deal with, since there is a possibility that the CPU cache will evict one of the cache lines without warning. The device will not observe this, and so might incur an access to main memory on the CPU node. In this paper we therefore focus on the case where the cache lines are homed on the device.
4.4. Discussion
A number of concerns might be raised about these protocols in practice; we address them in this section.
First is the issue of timeouts: we are blocking a core’s request for a cache line until a computation has completed. If this triggers a timeout in the CPU’s cache, the result is likely to be a memory fault exception, a machine check, or the processor simply locking up (on the Enzian, it is a machine check). We can work around these issues in practice by a small state machine on the device return a “not ready yet” response in the event of an approaching timeout, causing the software on the issue another request for the other cache line, extending the response time effectively indefinitely without the need for spinning. In practice, timeouts are quite long and the computations we target are short.
The second issue is more subtle, and concerns possible deadlock due to the device stalling a request until it can issue, and gets a reply back from, another to a different cache line. Such standards as exist for interoperable cache coherent interconnects are silent on what happens in this situation, possibly because it has not occurred until now. The design assumes that, in the platform, transactions on different cache lines can progress independently.
In most implementations, this assumption typically holds in order to maximize the memory bandwidth utilization and minimize request latency, and also to simplify reasoning about deadlock freedom of the coherence protocol itself. However, it may be the case that the cache stripes transactions across a limited number of independent units and which might deadlock if both A and B were mapped to the same unit. Were this to be the case, it could be avoided by careful placement of A and B in the physical address space. This is actually the case in our FPGA device implementation, described in Section 4.5.
4.5. Implementation in Enzian
We implemented both device-homed protocols in the Enzian platform, which we use for the evaluation in Section 5. The implementation uses SystemVerilog and exploits an existing “Directory Controller” on Enzian’s FPGA (Ramdas, 2024) which manages the coherent links to the CPU and handles transient states in the ECI protocol.
ECI provides all the messages we require to implement our messaging protocols. In particular, it includes the ability for the device to return a line to the CPU cache in Exclusive state, even though the CPU requested it in Shared. This optimization (discussed in Section 4.3) dramatically reduces latency, as shown in Section 5.1.
In our case, no flow-control issues arise since ECI uses separate virtual channels for request and reply messages, removing the chance of deadlock. However, other complications arise: ThunderX-1 generates Global Synchronization (GSync) messages which function as global barriers. A GSync blocks any subsequent transactions until all those started before the GSync have retired. But the invalidates sent are not awaited by the core they are still completed allowing the protocol to progress.
The ThunderX-1 divides its last-level cache functionality into a set of units called TADs, each of which can handle up to 16 simultaneous cache transactions. Two consecutive cache lines will be mapped to the different TADs by L2 to ensure that the operations can proceed independently.
On this CPU, barriers are needed to reliably order reads and writes. For example, it is critical that the core’s write buffer is drained into L1 cache (which is write-through to the L2) before a subsequent load from the core signals to the FPGA that the line can be pulled. The ARM architecture’s weak memory model requires a DMB barrier instruction between the writes and the read, and on the ThunderX-1 this is sufficient to ensure ordering observed at the FPGA.
On the FPGA, control (interception) is performed based on AXI read requests, since both Shared and Exclusive operations trigger these requests. The AXI writes go directly to the application. If the read is the result of the Exclusive, then the dummy data is sent back to the CPU as here is no fresh result yet.
5. Evaluation
In this section, we compare PIO over a coherent interconnect with two other reference points: PIO directly over PCIe– essentially, the CPU copying data to and from a region of PCIe address space – and descriptor-based DMA over PCIe. All use the same Enzian hardware platform. Coherent PIO uses the Enzian native ECI, while the two PCIe techniques connect the CPU and FPGA PCIe interfaces back-to-back with a loopback cable as described in Section 3, effectively turning the Enzian FPGA into a conventional PCIe FPGA card. DMA experiments all use the Xilinx XDMA hardware IP on the FPGA.
We explore three different application scenarios: synchronous RPC to an accelerator on the FPGA, a closely-coupled high-speed NIC, and a stream processor which moves some of its dataflow operator graph to the FPGA.
5.1. RPC to an accelerator
The first experiment measures null-RPC latency when the CPU invokes a function on the device, passing arguments, and obtaining a result. The protocol shown in 4(a) and detailed in Section 4.3 is used (i.e. the cache lines used for communication are homed on the FPGA).
Figure 5. Round trip ECI RPC latency, plus software latency between two sockets
Figure 5 shows the distribution of total latency for the this operation. We see a median latency of around 1600ns for the unoptimized protocol, wherein the result line is returned in Shared state as it was requested by the CPU. When we introduce the optimization of returning the line in Exclusive, the median latency drops to around 900ns.
For comparison, we also show an implementation of the FastForward protocol (Giacomoni et al., 2008a) exchanging cache lines between sockets in a dual-socket ThunderX-1-based Gigabyte R150-T61 server with similar CPU and DDR specification to Enzian; this achieves a median latency of about 1150ns.
The optimized protocol is performing 2 round-trip message exchanges over ECI, which has a one-way latency at the link later of about 150ns. The rest of the overhead (300ns) is mostly incurred in the protocol processing in the directory controller. The FPGA is clocked at about 300MHz.
A significant penalty is incurred when the results of the RPC are returned in Shared, resulting in further round trips over ECI to (unnecessarily) invalidate the line at the FPGA before sending the next message. This suggests that future coherent interconnects would benefit greatly from allowing this kind of operation.
Overall, the optimized protocol comfortably outperforms FastForward, despite the latter exploiting highly optimized hardware implementations at both ends (clocked at 2.1GHz) and aggressively polling, whereas our optimized protocol requires no spinning on the part of the CPU or device.
5.2. RPC over ECI vs. PCIe PIO and DMA
We now compare the RPC latency achieved using our protocol over ECI with PIO over PCIe and the Xilinx XDMA hardware using descriptors.
{listing}
[t] {minted} [ frame=lines, framesep=2mm, baselinestretch=1.2, bgcolor=LightGray, fontsize=, linenos, ] c // map 4KiB of PCIe memory for read/write int fd = open(”/sys/bus/pci/devices/…/resource0”, O_RDWR); void *pcie_bar = mmap(NULL, 4096, PROT_READ — PROT_WRITE, MAP_SHARED, fd, 0);
// write RPC arguments to PCIe memory *(volatile rpc_args_t *)pcie_bar = rpc_args; __sync_synchronize();
// read RPC result from PCIe memory rpc_res_t rpc_res = *(volatile rpc_res_t *)pcie_bar; __sync_synchronize(); Example RPC using PIO over PCIe on Linux. Error handling not shown for brevity.
For PIO over PCIe we pre-map the relevant PCIe aperture (BAR) into user space on the CPU and measure the time to write a value to this space and read a result back, as illustrated in Section 5.2. For XDMA numbers we use Xilinx XDMA benchmark (Xilinx DMA Benchmark, 2024), which essentially performs the same operation using descriptor-based DMA. In both these cases, the FPGA is configured to map the request and response into a write to and a read from onboard Block RAM, respectively.
As in Section 5.1, we measure latency to send a request from the CPU to the FPGA and read a response. However, this time we vary the payload size (of both request and response) and report median latency figures.
The results are shown in Figure 6. Up to payloads of 8KiB, the latency of XMDA-based communication is completely dominated by descriptor setup and manipulation, and remains largely independent of payload size. Faster PCIe interconnects *recall that this is a PCIe Gen3 x8 link) would only further emphasize this.
Latency over ECI is dramatically lower for smaller payloads. Unsurprisingly, it is constant up to 128 bytes (the size of a cache line on Enzian) and then increases linearly, since the protocol simply replicates each cache line exchange until all the data has been transferred. It is only at 8KiB that ECI RPC latency exceeds that of XMDA. One might argue that the transfer of multiple cache lines over ECI might be improved by clever pipelining and locating the lines in physical memory so as to exploit parallelism in the directory controllers, but we do not explore that in this paper.
Figure 6. RPC latency for different payload sizes
[…]
On this platform, using common techniques to improve performance like vector instructions has limited or no benefit for PIO over PCIe. For writes, the hardware already performs coalescing, and we see transactions of up to 512 bits regardless of vector size (confirmed using a logic analyzer). For reads, the 128 bit limit remains, and is responsible for the increase in latency between 16-byte and 32-byte payloads.
We conclude that up to the page size (4KiB) or the typical size of jumbo frames, PIO remains significantly faster than the DMA option commonly used for these transaction sizes. Thus even using the XDMA implementation for the experiments we still have the room for comparison with optimized DMA implementations.
5.3. Network interface
[…]
We implement 3 variations of a 100Gb/s NIC on Enzian. The first (7(a)) is a conventional approach connecting one of Enzian’s 100Gb Ethernet MACs fairly directly to the XDMA engines, resulting in a NIC which is programmed using descriptor rings. The second design (7(b)) buffers packets in onboard SRAM on the FPGA and allows the CPU to read and write packet and control data using PCIe reads and writes. The final approach (7(c)) uses a custom module on the FPGA to bridge between the ECI directory controller and the Ethernet MAC, using the protocol in 4(b) to deliver network packets to the CPU in cache lines. In all cases we clock the core NIC logic at 250 MHz.
(a) PCIe XDMA: packets transferred using descriptor-based DMA
(b) PCIe PIO: CPU reads descriptors and data from FPGA SRAM.
(c) ECI PIO: CPU interacts with PIO interface and SRAM through the L2 cache.
Figure 7. Different NIC architectures. Blue denotes data movers; red denotes memory that holds packet data.
For experiments we configure the Ethernet MAC in near-end PCS/PMA loopback mode for a reliable measurement of packet delivery inside the server without network delays. We define the receive latency as time passed from the last beat of the packet appearing on the ingress streaming interface of the Ethernet MAC, to the CPU fully receiving the packet content in its registers. The transmit latency is defined similarly as time passed from the CPU having the packet ready in registers, to the last beat of the packet appearing on the egress streaming interface on the MAC. We do not include the time it takes the packet to go through the Ethernet PCS/PMA loopback, since the MAC is fixed and the same in all implementations presented in this experiment.
We show figures for XDMA in both polled and interrupt-driven modes; in practice the difference in latency is small.
Figure 8 shows receive latency for different packet sizes for each of the NIC implementations. As in the RPC experiment, DMA latency is fairly stable across all packet sizes, suggesting that the latency is dominated by various DMA overheads, such as descriptor setup and manipulation.
Note that system calls here actually result in very little overhead (only a few microseconds); the overhead is actually dominated by cache misses manipulating DMA data structures in main memory.
[…]
Figure 9 shows transmit latency. In this case, PCIe PIO competes well with ECI across all packet sizes due to combining posted writes. PIO over ECI shows a slightly higher latency for small packets, due to it needing 2 round-trips over ECI for each 128 bytes, versus a single PCIe round trip. Both PIO solutions perform significantly better than DMA across all packet sizes.
These results show a clear advantage of using PIO over DMA, even for larger packets. The poor receive performance of PCIe-based PIO for receiving packets is largely due to a limitation of the PCIe design, that disappears for transmit.
5.4. Worked Example: Offloading Timely Dataflow
In this example we apply the techniques and results outlined so far to a practical workload: Timely Dataflow (Timely Dataflow, 2020), a real-time stream-processing system. We compare the baseline performance to a partially FPGA-offloaded implementation under three CPU–FPGA communication scenarios: DMA over PCIe, PIO over PCIe, and PIO over ECI.
Timely Dataflow schedules operators in a user-defined dataflow graph to evaluate complex operations on large streams of data, operating on variable-sized batches. Operators may be arbitrarily complex and are independent, and thus benefit from offloading where the overhead of shipping the data to the FPGA is compensated by greater throughput of the offloaded operators.
Figure 8. NIC implementations: receive latency
Figure 9. NIC implementations: transmit latency
To test the limits, we construct a worst-case scenario: a large (31-element) graph of simple operators (filters). The large graph maximizes communication overhead in the form of progress tracking data, while the simple operators leave the FPGA with only a small performance lead over the CPU. We want to predict whether PIO can make FPGA offload work in regimes where the fixed overheads of DMA make it impractical, for example where small batch sizes are needed to ensure data freshness (as in Differential Dataflow).
The first (blue) bars of Figure 10 show the baseline performance of the non-offloaded (CPU-only) system against batch size. Note that the horizontal axis is a log scale; All four curves are actually straight lines, with offsets and slopes obtained by a least-squares fit listed in Table 1 along with the coefficient of determination ($r^{2}$). The remaining bars cover the three offload cases. Each is the median of 1000 samples.
While no offload configuration beats the baseline, the two PIO cases outperform DMA up to batch sizes of 512B (for PCIe) and 4KiB (for ECI), with ECI almost matching the baseline at the smallest batch size (128B). To understand why offloading fails to produce a speedup in this example, and what changes will be necessary to allow it to, we look more closely at the fitted coefficients in Table 1.
As shown by the $r^{2}$ values, the fit for each dataset is excellent. We can thus model the time taken to process a batch as a fixed component, or overhead, independent of the amount of data processed (e.g. running the Timely scheduler, or setting up DMA descriptors) plus a component which scales linearly with the size of the batch.
Compared to the baseline, DMA offload has a significantly lower per-byte cost ($3.7$ vs. $8.7ns/B$) but higher fixed cost ($336\mu s$ vs. $204\mu s$). This includes the cost of descriptor setup and interrupt handling, while the per-byte cost shows the higher throughput of the FPGA. This includes both computation and data transfer but as DMA makes very efficient use of the (PCIe 3.0 x8) interconnect, once setup costs are paid, the per-byte transfer cost is on the order of $0.1ns$, allowing us to estimate the per-byte cost of FPGA compute at $3.6ns/B$. The CPU figure represents only compute.
This linear model predicts a break-even at 26KiB. This matches the observed data (not plotted) where DMA offload first outperforms the baseline at a batch size of 32KiB. Wherever the FPGA compute cost is at least $0.1ns$ lower per byte than the CPU, DMA offload will outperform for large enough batches. However, due to the large fixed overhead it can never outperform the CPU for batches smaller than 16KiB. This limits its usefulness where small batch sizes are needed.
While they make less efficient use of the interconnect bandwidth (as indicated by their higher per-byte costs), their lower fixed overhead allows the PIO modes to significantly outperform DMA at smaller batch sizes. What would it take for the faster of the two (ECI) to outperform the CPU at a moderate batch size?
As the FPGA pipeline is identical in all cases, the limiting factor is the per-byte transfer cost. Taking our above estimate of FPGA compute at $3.6ns/B$, we can estimate this cost at $30-3.6=26.4ns/B$. Letting $S$ stand for batch size and $C$ for this transfer cost, the two break even where
$204\mu s+S\times 8.7ns/B=221\mu s+S\times(C+3.6ns/B)$
Rearranging gives
$C=5.1ns/B-17\mu s/S$
5.1ns/B being the speedup due to FPGA processing and $17\mu s$ the fixed overhead relative the baseline, the negative sign due to the higher-than-baseline overhead.
From this we see that given the overhead the smallest batch for which a speedup is possible is around 3300B, and that a batch size of 4096B demands a transfer cost not more than 0.95ns/B, or 120ns per 128B cache line. As the existing protocol serializes on each line, this cost includes the round-trip latency. As the hardware can handle many in-flight transfers at once, modifying the protocol to transfer 32 lines at once would reduce the per-byte cost enough that ECI-PIO offload breaks even.
Alternatively or additionally, bringing the fixed cost down reduces the smallest practical batch. If the relative overhead could be brought below 500ns/batch then with parallel cache lines giving a 1ns/B transfer cost, FPGA offload using ECI-PIO would outperform the CPU from a batch size of 128B, or a single cache line. With such an optimized protocol, ECI-PIO would outperform all other cases until a batch size of 142KiB, above which PCIe-DMA would overtake it giving a net offload benefit across all batch sizes.
Figure 10. DMA and PIO performance for Timely Dataflow FPGA offload with a graph size 31
time/byte ($ns$) overhead ($\mu s$) $r^{2}$
Baseline 8.7 204 1.00
PCIe-DMA 3.7 336 0.99
PCIe-PIO 88 277 1.00
ECI-PIO 30 221 1.00
Table 1. 1st-order (offset + slope) coefficients fitted to the data series of Figure 10, as cost per byte and fixed overhead
6. Related work
6.1. Coherent interconnects
Coherent replacements or extensions of existing device interconnects have been under development for some time and are beginning to see widespread availability. IBM’s OpenCAPI (Stuecheli et al., 2018) builds on PCIe by adding a protocol layer for coherence. It and the competing Gen-Z (Keeton, 2015) and CCIX (CCIX Consortium and others, 2024) protocols have meanwhile been either merged into or replaced by CXL (Li et al., 2023), which seems likely to be the standard interoperable protocol in the short- to medium-term future.
Commercial CXL hardware implementing the more recent revisions, which allow less restricted use cases, is still not widely available. Consequently research on the software uses of coherent interconnects is still largely performed on specifically research-oriented machines such as the Intel HARP (Intel, 2024; Intel Harp, 2024), and Enzian (Cock et al., 2022). While HARP provides a pre-configured coherent cache in the FPGA with a conventional interface, Enzian’s ECI protocol is fully open and customizable.
Other notable coherent interconnects include the RISC-V-specific TileLink (Berkeley Architecture Research, 2022; Terpstra, 2017; Cook et al., 2017) and NVIDIA’s NVLink 2.0 (Nvidia, 2024; Wei et al., 2023a). TileLink is an open standard and suitable for research but hasn’t yet been implemented in server-scale hardware. NVLink is proprietary and closed.
Existing work has built on the available research platforms to explore the design space for practical applications. Centaur (Owaida et al., 2017) demonstrated the FPGA offload of database operations on Intel HARP v1, and Cohort (Wei et al., 2023b; Mukherjee et al., 1996) proposed and prototyped a unified queue-based software interface to such accelerators building on the P-Mesh interconnect of OpenPiton (OpenPiton Open Source Research Processor, 2024).
Open, extensible protocols have been used to explore coherent offload in simulation, including making the case for protocol specialization building on the Spandex protocol family (Alsop et al., 2022, 2018). Conda (Boroumand et al., 2019) likewise employed simulation to explore the design space for coherent device interconnects, with a focus on reducing unnecessary message traffic.
The Denovo protocol (Sinclair et al., 2015) is presented as an improvement specifically for CPU-GPU coherence, tailoring the protocol to match the comparatively predictable access patterns of typical GPU workloads.
6.2. Cache line-based communication
Other work (Tamimi et al., 2022) has already noted that the cache line is a better unit of transfer where small operations are expected. This is a common observation, long since applied in the original FastForward (Giacomoni et al., 2008b) and Barrelfish UMP (Baumann et al., 2009b) protocols. More recently Concord (Iyer et al., 2023) communicates scheduling decisions between workers and the dispatcher via a polled cache line, converting worker threads from interrupt-driven to poll-mode “CPU drivers”. A similar technique is used in Shinjuku (Kaffes et al., 2019).
A recent study of data center RPC from Google (Karandikar et al., 2021) reinforces the importance of small transfers and highlights the large latency of PCIe transactions.
6.3. Exploring PIO
The space where PIO is preferable to DMA is likewise well studied. The hXDP (Brunella et al., 2020) FPGA NIC work highlighted the high overhead of small PCIe transactions, and consequently performs small batch computations solely on the CPU.
In kPIO+WC (Larsen and Ben Lee, 2015) the authors demonstrate the benefit of write combining for PCIe-based PIO. Compared to a traditional DMA NIC, their FPGA-based prototype shows better latency and throughput for small and medium-sized messages and comparable throughput for large messages. The ThunderX CPU automatically performs combining for PCIe writes, and the results of Section 5.2 thus reflect this optimization.
Neugebauer et. al. (Neugebauer et al., 2018) provide a comprehensive analysis of PCIe performance in the context of NICs.
Dagger (Lazarev et al., 2021) builds on CCI-P, the commercialized implementation of HARP’s coherent cache, to construct an FPGA-based NIC specialized for low latency RPC. A host-coherent cache holds connection states and the necessary structures for the transport layer on the NIC, while the payload remains in host memory. This minimizes FPGA memory demands, and exploits the lower latency of cache misses compared to conventional PCIe-based NICs.
7. Conclusion and Future Work
In this paper we have demonstrated that PIO on modern hardware outperforms DMA over a wide range of payload sizes in a number of different applications. For anything up to a disk block, VM page, or Ethernet Jumbo Frame, PIO offers (often dramatically) lower transfer times mostly thanks to its low fixed overheads. This advantage is consistent across RPC workloads, NIC packet transfer, and FPGA offloaded applications.
This strongly contradicts the accepted wisdom, under which these specific payloads and applications are the canonical examples of DMA-suitable tasks. Moving beyond conventional interconnects (PCIe), we have shown that the advantage grows even greater for the emerging class of coherent interconnects. This is partly due to a more suitably-sized unit of transfer (the cache line) and partly due to the opportunity to interact with the coherence protocol in a carefully-optimized manner.
Moving back to PIO makes sense in a landscape of plentiful, cheap CPU cores and fine-grained RPC-style workloads.
This work is ongoing, and we intend to develop it further in several ways. Both the cache-coherent NIC and Timely Dataflow offload designs will continue to be optimized, to explore the limits of PIO-style device communication on emerging coherent interconnects. We will also further explore the design of appropriate programming interfaces and OS-level integration of coherent mechanisms for high-performance devices.

### v2

@@
 In this paper we question this wisdom in the light of modern hardware and workloads, particularly in cloud servers.
-We argue that the assumptions that led to this model are obsolete, and in many use-cases use of programmed I/O, where the CPU explicitly transfers data and control information to and from a device via loads and stores, actually results in a more efficient system.
-We quantitatively demonstrate these advantages using three use-cases: fine-grained RPC-style invocation of functions on an accelerator, offloading of operators in a streaming dataflow engine, and a network interface targeting for serverless functions.
-Moreover, we show that while these advantages are significant over a modern PCI Express (PCIe) peripheral bus, a truly cache-coherent interconnect offers significant additional efficiency gains.
+Like others before us, we argue that the assumptions that led to this model are obsolete, and in many use-cases use of Programmed I/O (PIO), where the CPU explicitly transfers data and control information to and from a device via loads and stores, actually results in a more efficient system.
+However, unlike others to date, we push this idea further and show, in a real implementation, the gains in average and tail latency for fine-grained communication achievable using an open cache-coherence protocol which exposes cache transitions to a smart device.
+We show this using three use-cases: fine-grained RPC-style invocation of functions on an accelerator, offloading of operators in a streaming dataflow engine, and a network interface targeting for serverless functions, comparing our use of coherence with both traditional DMA-style interaction and a highly-optimized implementation using PIO over PCI Express (PCIe).
 1. Introduction
-There is a current consensus on how CPUs should communicate with modern, high-performance devices like network interface adaptors (NICs), storage controllers, Graphics Processing Units (GPUs), and other computational accelerators: for the most part, this is achieved using DMA and rings of descriptors to decouple CPU and device as much as possible.
-On modern PCIe interconnects, this consensus results in impressive bandwidth for applications which transfer data in large chunks.
-However, in this paper we question this consensus in the light of several factors: (1) important workloads involving irregular access and fine-grained interaction between CPU and device, (2) increasing core counts in servers, (3) data center applications where many machines are executing the same application on most of their cores, and (4) the emergence of cache-coherent peripheral interconnects, particularly those like CXL.cache 3.0 (CXL Consortium, 2024a) in which devices are first-class participants in the coherence protocol.
-We argue that these factors make Programmed I/O (PIO), where CPU software performs I/O by directly reading and writing device registers, a better mechanism for data transfer in some important scenarios than one based on DMA.
-We use a real hardware platform which implements a coherent interconnect between a server-class CPU and a large FPGA to investigate the trade-offs between DMA using descriptor rings, PIO directly over a PCIe interconnect, and PIO over a cache-coherence protocol.
+Modern interfaces between CPUs and high-performance devices like network interface adaptors (NICs), storage controllers, Graphics Processing Units (GPUs), and other computational accelerators are designed to optimize throughput.
+Built over a PCIe interconnect, descriptor rings in main memory hold queues of operations and completions, and the device performs DMA transactions to read and write both descriptors and payload data.
+This design, along with the underlying PCIe interconnect, trades off latency for small transactions in favor of high throughput for large ones.
+In this paper we question this consensus in the light of workloads which depend for performance not on bulk throughput, but the cumulative latency of small transactions between CPU and device: closely-coupled hardware accelerators for irregular workloads, or data center Remote Procedure Call (RPC) using small messages.
+We do this in the context of emerging cache-coherent peripheral interconnects like CXL.cache 3.0 (CXL Consortium, 2024).
+The continuing increase in PCIe bandwidth means that, for a given transfer size, the significance of PCIe latency (time to first byte) has become more of a concern, leading to many proposals for reducing the overhead of descriptor management (which we survey in Section 7.
+However, we propose a more radical approach, aimed at use-cases where reduced latency is more important than higher maximum throughput: using PIO to Memory-Mapped I/O (MMIO) devices directly, and avoiding DMA, descriptor queues, operation buffering, and interrupts entirely.
+We use a real hardware platform which implements a coherent interconnect between a server-class CPU and a large FPGA to investigate the trade-offs between DMA using descriptor rings, PIO directly over a PCIe interconnect, and PIO over a full cache-coherence protocol.
 We show that, for small (around 1KiB or less) transfers, PIO posted writes over PCIe significantly outperform DMA, although reads are less efficient since, unlike writes, they cannot be pipelined and so incur PCIe’s significant round-trip latency penalty.
 However, a full cache coherence protocol which avoids PCIe significantly outperforms both DMA and PIO over PCIe.
+Crucially, not only is median latency considerably lower, but the tail of the latency curve is reduced significantly when using PIO over PCIe, and completely eliminated using the cache coherence protocol.
 We go further, however, and show how a conventional MESI-like coherence protocol can be exploited in novel ways by intelligent devices which can interact with an unmodified CPU at the coherence message level.
@@
 In Section 3 we describe and motivate the experimental hardware platform we use for our evaluation, and also provide calibrating performance comparisons with PC data center servers.
+We show how to use PIO with PCIe in Section 4.
 In Section 5, we discuss why true cache-coherent device interconnects are fundamentally different from today’s PCIe platforms, and present a set of efficient protocols for passing messages with low latency between software running on an unmodified CPU, and a smart device which has low-level access to the coherence protocol.
-In Section 4.5, we go on to present the implementation of some of these protocols on our experimental platform.
+In Section 5, we go on to present the implementation of some of these protocols on our experimental platform.
 Section 6 then compares a traditional DMA-based approach with PIO over PCIe and our new coherence-based protocols using three different use-cases: (1) lightweight, synchronous local invocation of functionality on a computational accelerator, (2) I/O to and from an intelligent network interface, and (3) hardware offload of a distributed dataflow stream processor to an FPGA.
@@
 Our work in this paper is motivated by the interaction between modern trends in platform interconnect and also the changing ways in which these interconnects are used, particularly in data center and cloud servers.
-2.1. Background: hardware/software interaction
+Interconnects and devices:
 In early machines, the CPU interacted with peripheral devices exclusively using programmed I/O (PIO): the device exposed hardware registers mapped into the processor’s address space.
 Both data transfer and control of the device involved reads and writes by the CPU to these registers.
+This model is also known as memory-mapped I/O (MMIO).
 This model made sense when devices were extremely simple, particularly when augmented by the use of device interrupts to obviate the need for the CPU to poll status registers on the device waiting for an I/O operation to complete.
@@
 This model is optimized for throughput – it is most efficient when a large volume of data must be transferred.
-2.2. Trends in interconnects and devices
+On modern PCIe interconnects, this consensus results in impressive bandwidth for applications which transfer data in large chunks, up to 64GiB/s for PCIe 5.0 x16.
 […]
-Recent trends in accelerator design, however, have led to proposed standards for cache coherent peripheral interconnects, and the related idea of extending processor cache coherence to heterogeneous devices: accelerators, networking interfaces, and storage controllers (Stuecheli et al., 2018; CCIX Consortium and others, 2024; CXL Consortium, 2024b; Nvidia, 2024; AMBA® CHI Architecture Specification, 2024).
+For the large batch transfers used in GPU-derived programming models (e.g., machine learning workloads), this latency penalty is insignificant when amortized over a large transfer.
+Indeed, such software has arguably adapted to this model imposed by the hardware (Hooker, 2021).
+Recent trends in accelerator design, however, have led to proposed standards for cache coherent peripheral interconnects, and the related idea of extending processor cache coherence to heterogeneous devices: accelerators, networking interfaces, and storage controllers (Stuecheli et al., 2018; CCIX Consortium and others, 2024; CXL Consortium, 2024; NVIDIA, 2024; AMBA® CHI Architecture Specification, 2024).
 The more far-reaching of these proposals are fully symmetric, i.e.
 devices are first-class participants in a distributed directory-based cache coherence protocol.
-Communication in software between peer modes in a cache-coherent system is typically very different from using descriptor-based DMA.
-FastForward (Giacomoni et al., 2008a), Barrelfish’ UMP (Baumann et al., 2009a), Concord (Iyer et al., 2023), and Shinjuku (Kaffes et al., 2019) adopt a much more direct approach to sending small messages with low latency, e.g.
-by exploiting the cache coherence protocol to transfer lines between caches on demand, and local polling to provide synchronization.
-Recently, cache coherence between NUMA CPUs has been used to simulate communicating between software and a hypothetical cache-coherent NIC (Schuh et al., 2024).
-2.3. “New” workloads
-Without denying the importance of throughput-oriented data transfer between devices and cores for workloads like current machine learning training and inference, it is worth stepping back and questioning many of the assumptions that led us here, and what other workloads might be a poor fit today for this model.
-Any workload involving fine-grained, frequent interaction between the CPU and another device incurs a significant latency penalty when using DMA for data transfer.
+Fine-grained workloads:
+While throughput-oriented workloads involving data transfer between devices and cores (e.g., AI training and inference) are hugely important today, we step back in this paper and consider other workloads that might be a poor fit today for this model, because they involve fine-grained, frequent interaction between the CPU and another device.
 Many irregular workloads have this property (Gioiosa et al., 2017; Bean, 2016; Kashimata et al., 2019; Zeng and Prasanna, 2020).
-Moreover, data center messaging (essentially, network Remote Procedure Calls (RPCs)) exchange many small messages (Karandikar et al., 2021; Lazarev et al., 2021).
-While end-to-end latency is mostly network propagation time in this case, latency incurred in the end system is a good proxy for resources (such as CPU cycles) wasted before invoking the application code, and in marshaling and sending any reply.
-There are also several assumptions underlying the DMA-based communication model which are worth revisiting.
-In particular, CPU execution and I/O were decoupled because CPU cycles were precious (there being few cores), and the CPU had plenty of different tasks to work on at any time.
+Moreover, data center messaging (essentially, RPCs) exchange many small messages (Karandikar et al., 2021; Lazarev et al., 2021).
+While end-to-end latency is mostly network propagation time in this case, latency incurred in the end system is a good proxy for key resources (such as CPU cycles) wasted before invoking the application code, and in marshaling and sending any reply.
+All these workloads incur a significant latency penalty when using DMA for data transfer. Moreover, the assumptions which led to the common DMA descriptor-based model do not hold for these workloads.
+For example, CPU execution and I/O were decoupled using descriptor rings because CPU cycles were precious (there being few cores), and the CPU had plenty of different tasks to work on at any time.
 This is no longer true.
-Server CPUs have had many cores for some time, and moreover when running data center applications they are often dedicated to a single application, sometimes with miscellaneous functions moved to a small subset of the cores (Marty et al., 2019).
-Another example: for throughput-oriented workloads DMA has evolved to efficiently transfer data to and from main memory without polluting the CPU cache.
-However, for small message applications, it is important that almost all the data gets into the right CPU cache as quickly as possible.
-This leads us to rethink the use of the DMA model, and whether directly involving the CPU in data transfers to and from devices might be a better approach for many scenarios on modern hardware.
-Others have also suggested this in some cases (Larsen and Ben Lee, 2015; Chaudhari et al., 2017), but we also push the question much further: given a future cache-coherent interconnect, what new ways of using Programmed I/O might become compelling for data center applications.
+Server CPUs now have 10s or 100s of cores, and moreover when running data center applications they are often dedicated to a single application, sometimes with miscellaneous functions moved to a small subset of the cores (Marty et al., 2019).
+Most of the CPU cores have nothing to do except handle small messages coming from the network.
+As a second example: for throughput-oriented workloads DMA has evolved to efficiently transfer data to and from main memory without polluting the CPU cache.
+However, for small, fine-grained interactions, it is important that almost all the data gets into the right CPU cache as quickly as possible.
+These reasons led us to rethink the DMA-based model, and whether directly involving the CPU in data transfers to and from devices might be a better approach for these scenarios on modern hardware.
+We are not the first to suggest this direction for some cases (Larsen and Ben Lee, 2015; Chaudhari et al., 2017), but we push the question much further: given a future cache-coherent interconnect, what new ways of using Programmed I/O might become compelling for data center applications?
+Note that we are not proposing abandoning the traditional model, or some of its enhancements we survey in Section 7, for throughput-oriented applications, which can continue to ride the current hardware curve for bandwidth.
+Instead, we argue it is time to examine complementary alternatives.
 3. Experimental platform
-We investigate the modern trade-offs between PIO, DMA, and cache-coherent interconnects using the Enzian research computer (Cock et al., 2022).
-Enzian is useful as a placeholder for future, interoperable, cache-coherent interconnects like CXL.
-We use a real hardware platform rather than a simulator.
+Figure 1. XDMA invocation latency comparison.
+Figure 2. PIO invocation latency comparison.
+We investigate the modern trade-offs between PIO, DMA, and cache-coherent interconnects using a real hardware platform rather than a simulator.
 Simulation would be appropriate for detailed quantitative comparisons of known techniques, but in this paper we explore unconventional uses of cache coherence and so want to establish not only that our techniques are performant, but that they can practically be implemented in a real system.
-An Enzian machine can be thought of as a two-socket NUMA server, where one socket houses a Marvel Cavium ThunderX-1 CPU, with a Xilinx XCVU9P FPGA in the other.
+The Enzian research computer (Cock et al., 2022) can be thought of as a two-socket NUMA server, where one socket houses a Marvel Cavium ThunderX-1 CPU, with a Xilinx XCVU9P Field Programmable Gate Array (FPGA) in the other.
 The CPU is a 48-core ARMv8 processor running at 2.0GHz, and 128GiB of 2133MT/s DDR4 memory spread across 4 memory controllers.
@@
 The CPU and FPGA are connected by the CPU’s native inter-socket cache coherent protocol, the Enzian Coherence Interface (ECI), which Enzian implements on the FPGA.
-It is a MOESI-like directory based protocol, and the physical address space is statically partitioned between the two NUMA nodes (CPU and FPGA).
+It is a MOESI-like directory protocol with the physical address space statically partitioned between the NUMA nodes (CPU and FPGA).
 It uses 24 bidirectional lanes (organized into 2 12-lane links), each running at about 10Gb/s, for a theoretical inter-socket bandwidth of about 30GiB/s.
@@
 In this work we connect these two interfaces with a loopback cable; this essentially replicates the case of an PCIe accelerator card and allows us to compare PCIe with ECI in our experiments.
-3.1. Characterizing the platform
-Enzian provides a real hardware platform with server-class performance to evaluate different forms of device communication.
-However, direct comparisons with modern server platforms is not straightforward: the latter do not offer a full cache coherence protocol between the CPU and devices, but do provide more recent and therefore higher-performance cores, memory systems, and PCIe interconnects.
-Therefore, in order to calibrate against a modern PC for both PIO and DMA over PCIe, we present two benchmark comparisons of Enzian with a modern PC (Intel Core i7-7700 3.6GHz Kaby Lake, PCIe Gen.3 x16) with cache line size of 64 bits connected to an AMD Xilinx Virtex UltraScale+ VCU118 card, using the same “write then read” experiment detailed in Section 5.1.
+While Enzian provides a real hardware platform with server-class performance, direct comparison with modern server platforms is not straightforward: the latter do not offer full cache coherence between CPU and devices, but have more recent, faster cores, memory systems, and PCIe.
+We therefore calibrate against a modern PC for both PIO and DMA over PCIe with benchmark comparisons of Enzian and a modern PC (Intel Core i7-7700 3.6GHz Kaby Lake, PCIe Gen.3 x16) with 64-byte cache lines connected to an AMD Virtex UltraScale+ VCU118 card, using the “write then read” experiment detailed in Section 6.1.
 The VCU118 uses the same FPGA as Enzian, albeit a slightly slower speed gauge.
-3.1.1. DMA performance over PCIe
-The first experiment uses DMA, using the Xilinx XDMA IP and associated descriptor-based protocol to transfer data between the CPU and the FPGA over PCIe.
+DMA performance over PCIe:
+The first experiment uses the Xilinx XDMA IP and its descriptor-based protocol to transfer data between CPU and FPGA over PCIe.
 Figure 1 shows results for various data transfer sizes, running the Xilinx driver in both interrupt-driven and polled modes.
-Figure 1. Comparison of XDMA latency for RPC workload between Enzian and PC
-We observe that, using PCIe, exchanging messages using XDMA is about 3 times as fast on the PC server than on Enzian, while the difference between interrupt-driven and polling performance is much less significant.
-The latency of a single transaction is pretty constant on both platforms up to the PCIe transaction size limit of 4KiB, and then increases.
-The difference in performance between the two architectures is due to several factors.
-For very small transfers, the CPU overhead of setting up descriptors dominates, and an x86 core is simply faster than the slower (and mostly in-order) ThunderX-1 core.
+We see that XDMA over PCIe is about 3 times faster on the PC than on Enzian, while the difference between interrupt-driven and polling performance is much insignificant.
+Single transaction latency is almost constant on both platforms up to the PCIe transaction size limit of 4KiB, and then increases.
+The performance difference between the machines is due to several factors.
+For small transfers, the CPU overhead of descriptor setup dominates, and an x86 core is simply faster than a ThunderX-1 core.
 In addition, the memory system is somewhat faster on the PC.
 The factor of 2 difference in PCIe bandwidth between the two platforms does not appear to be a factor in these experiments.
-3.1.2. PIO performance over PCIe
-Our second calibration experiment involves the CPU itself doing PIO reads and writes over PCIe to the FPGA.
-We experimented with several different optimizations in this experiment, in particular the use of write combining and vector instructions.
-We present the best figures we could obtain on both platforms.
-Figure 2. Comparison of PIO latency for RPC workload between Enzian and PC
-Figure 2 shows the results.
-For transaction sizes above 32 bytes, the PC is about twice as fast as Enzian over the PCIe cable between CPU and FPGA.
-Here, the factor of 2 difference in PCIe bandwidth is responsible for the latency difference, in particular for reads (which are limited to 128 bits on both platforms).
-Writes, in contrast, are combined and highly pipelined at the PCIe interface.
-4. Programmed I/O over coherent interconnects
-In this section we explore what programmed I/O might look over coherent interconnects.
-It is of course possible for a CPU to simply copy data over a coherent interconnect as in conventional programmed I/O.
-This requires careful management of the processor’s cache hierarchy with explicit cache flush and invalidate instructions, but can be appropriate when the underlying hardware standard itself mandates a cache on the device, as with OpenCAPI 4.0 (Stuecheli et al., 2018; Allison et al., 2020).
-That said, here we are more interested in how a coherent interconnect can best be used for low-latency, efficient movement of data between CPU registers and a device, including cases where the conventional notion of the cache as a view over memory contents is not appropriate (such as data arriving over a network).
-We scope the question further as follows: we consider the case where the CPU cores, or their Last-Level Cache (LLC), communicate with devices via a MESI-like (Nagarajan et al., 2020), message-based coherence protocol with stable states corresponding to Modified (dirty), Shared (clean), Exclusive (clean), and Invalid, using a distributed cache directory.
+PIO performance over PCIe
+In the second calibration experiment the CPU performs PIO reads and writes over PCIe to the FPGA.
+We experimented with several different optimizations in this experiment (see Section 4), in particular the use of write combining and vector instructions.
+We present the best figures obtainable on each platform.
+Figure 2 shows that for transaction sizes above 32 bytes, the PC is about twice as fast as Enzian over the PCIe cable between CPU and FPGA.
+Here, the difference in PCIe bandwidth is responsible for the latency difference, in particular for reads (limited to 128 bits on both platforms).
+Writes, in contrast, are combined and pipelined at the PCIe interface.
+4. Programmed I/O over PCIe
+PCIe is complex, and we spent considerable time optimizing CPU-initiated reads and writes for different platforms.
+We describe here how we arrived at optimally-efficient PIO latency to the FPGA over PCIe.
+We pre-map PCIe apertures (BARs) into user space on the CPU and measure latency as the time to write a value to this space (thereby invoking a function on the device) and read a result back (Figure 3).
+// map 4KiB of PCIe memory for read/write
+int fd = open("/sys/bus/pci/devices/.../resource0",
+    O_RDWR);
+void *pcie_bar = mmap(NULL, 4096,
+    PROT_READ | PROT_WRITE, MAP_SHARED, fd, 0);
+// write arguments to PCIe memory
+*(volatile invoke_args_t *)pcie_bar = invoke_args;
+__sync_synchronize();
+// read result from PCIe memory
+invoke_res_t invoke_res = *(volatile invoke_res_t *)pcie_bar;
+__sync_synchronize();
+Figure 3. An invocation using PIO & PCIe. Error handling omitted.
+PCIe memory writes are posted transactions and may complete out-of-order, allowing multiple requests in-flight at the same time.
+Most systems also have dedicated bus units to combine writes and coalesce transactions.
+Enzian’s ThunderX-1 performs write-combining for stores to PCIe, allowing 512 bits per bus round-trip (confirmed with a logic analyzer) and resulting in efficient PIO write transactions.
+[…]
+This means that on ThunderX-1, common optimization techniques like vector instructions have negligible benefit for PIO over PCIe: the hardware already coalesces writes into 512-bit transactions and reads are limited by the read bus width of 128 bits.
+x86 has more optimization opportunities for PIO reads and writes.
+The Intel Data Streaming Accelerator (Intel, 2024) offers MOVDIR64B and ENQCMD instructions, allowing 64-byte PIO writes with write-combining mappings and user-space descriptor submission without driver intervention.
+Mapping memory with write-through attributes, as suggested by Tide (Humphries et al., 2024), allows cache line-sized PIO reads.
+These architecture-specific extensions are not available on ThunderX-1.
+5. PIO over a coherent interconnect
+We now turn to programmed I/O over a cache coherent interconnect.
+A naive approach would have the CPU simply copy data over a coherent interconnect as in conventional PIO.
+This requires careful cache management on the CPU using explicit cache flush and invalidate instructions, but is appropriate when the hardware standard mandates a cache on the device, as with OpenCAPI 4.0 (Stuecheli et al., 2018; Allison et al., 2020).
+However, we are more interested in how a coherent interconnect can best perform low-latency, efficient data movement between CPU registers and device hardware, including cases where the conventional notion of the cache as a view over memory contents is not appropriate, such as data arriving over a network.
+Scoping:
+We scope the scenario further to the case where the CPU cores, or their Last-Level Cache (LLC), communicate with devices via a MESI-like (Nagarajan et al., 2020), message-based coherence protocol with stable states corresponding to Modified (dirty), Shared (clean), Exclusive (clean), and Invalid, using a distributed cache directory.
 Note that this means that each address or cache line has a home node, which maintains the directory entry for the line.
-Moreover, we assume a symmetric protocol: each device is a first-class participant in the protocol with a subset of physical addresses homed at it, and the device therefore maintains directory state for these lines.
-While these assumptions are rare in hardware today, they are true both of the proposed CXL.cache standard (CXL Consortium, 2024a) and the Enzian platform we use for experiments.
-4.1. Baseline: naively communicating with coherence
-Our starting point is existing techniques which use cache coherence to perform inter-core communication, for example  (Giacomoni et al., 2008a) or Barrelfish UMP  (Baumann et al., 2009a).
-The basic idea is shown in Figure 3; to software, the sender writes data to a cache line on which the receiver is spinning (reading in a tight loop).
-Initially, the line is in (say) Shared state in both caches.
-When the sender writes the line (typically in one go due to the core’s write-combining buffer), it transitions briefly to Modified state, before being fetched by the receiver over the interconnect and reverting to Shared.
+We also assume a symmetric protocol: each device is a first-class participant in the protocol with a subset of physical addresses homed at it, and the device therefore maintains directory state for these lines (although it might not actually cache data itself).
+Moreover, we take the perspective of the device designer: the device sees individual cache messages, rather than transparently coherent memory.
+These assumptions are rare in hardware today but they do hold for proposed standards like CXL 3.0 (CXL Consortium, 2024), and also for the Enzian platform we use in experiments.
+Baseline: simple coherence
+We first discuss using simple cache coherence to re-apply existing techniques inter-core communication in software, such as FastForward (Giacomoni et al., 2008).
+Here (Figure 4), the sender (CPU) writes data to a cache line on which the receiver is spinning reading.
+Assume the line starts in Shared state in both caches.
+When the sender writes the line (typically atomically via the core’s write buffer), it transitions briefly to Modified state, before being fetched by the receiver over the interconnect and reverting to Shared.
 Figure 4. Simple inter-core communication over coherence
-This can be highly efficient in software, particularly when deployed as a ring buffer of cache lines, and results in very little interconnect traffic (since spinning is usually on a local copy of the line).
-One can easily replace one core with a device in this scheme (with suitable state machine) to obtain an I/O protocol.
-The downside is, of course, the CPU cost of spinning, both in wasted cycles and the inability to switch to another task.
-The design of protocols like FastForward is based on the idea that each endpoint can only manipulate its local cache using software instructions: loads, stores, etc.
-Moreover, it is assumed (reasonably) that flush or invalidate instructions should be avoided because they might be slow, and might require kernel privilege.
+This can be highly efficient in software, particularly when deployed as a ring buffer of cache lines, and results in minimal interconnect traffic (since the receiver spins on a local copy of the line), although the CPU itself must spin.
+Indeed, this is the technique recently used in CC-NIC (Schuh et al., 2024) to emulate a NIC using a different CPU socket.
+However, this design assumes that each endpoint can only manipulate its local cache using software instructions (loads, stores, etc.) and also that flush or invalidate instructions should be avoided for performance reasons.
 This is not the case when one of the endpoints is a device with message-level access to the cache coherence protocol, and possibly maintaining its own directory state.
-4.2. The implications of low-level access
-Unlike software running on a core behind a cache, a device participating in a coherence protocol has visibility over more state and events, and can also initiate more operations.
-Note that there is no need for a cache per se on the device at all, indeed this is usually inappropriate.
-First, the device receives messages from the CPU cache that request lines in Shared or Exclusive state, or which request that the caching state on the be downgraded (e.g.
-from Exclusive to Shared, or from Shared to Invalid).
-It is therefore able to react internally to such events as it chooses, as long as it does not violate the protocol.
+The implications of message-level access
+A device on a symmetric coherent interconnect can see more cache state and events, and also initiate more operations, than a CPU core can.
+Note that this does not entail a cache per se on the device at all, indeed this is rarely appropriate.
+First, the device receives messages from the CPU cache that request lines in Shared or Exclusive state, or which request that the caching state on the device be downgraded (e.g., from Exclusive to Shared, or from Shared to Invalid).
+It is free to react internally to these events as it chooses, as long as it does not violate the protocol.
 Second, the device can issue such requests itself at any time, allowing more fine-grained control of the state of a line in the CPU’s cache.
@@
 Instead, it can choose to delay the response (blocking the requesting core) until some other operations complete.
-Care must be taken to respond before any hardware-imposed timeouts, but these are typically generous (hundreds of milliseconds on the ThunderX-1, for example) and can be worked around in software (e.g.
-by sending a “try again” response within the timeout).
+Care must be taken to respond before any hardware-imposed timeouts, but these are typically generous (hundreds of milliseconds on the ThunderX-1, for example) and can be worked around in software (e.g., by sending a “try again” response within the timeout).
 Finally, the device can choose to interpret events like remote requests as more than simply loads or stores to the line concerned.
@@
 Figure 5. Protocol variants for efficient CPU-device messaging
-4.3. Communication protocol variants
-We present 4 different but similar communication protocols (shown in Figure 4) which vary based on (a) the CPU or the device initiates the message transfer, and (b) the cache lines used are homed on the CPU or device.
-All provide a synchronous RPC: a transfer of a single cache-line-sized message, followed by a single cache-line response.
+Communication protocol design variants
+We present 4 different but similar communication protocols (shown in Figure 5) which vary based on whether (a) the CPU or the device initiates the message transfer, and (b) the cache lines used are homed on the CPU or device.
+All provide a synchronous invocation: a transfer of a single cache-line-sized message, followed by a single cache-line response.
 All use 2 cache lines, which we designate A and B.
@@
 We focus on the simple 2-line case in this paper.
-The protocols use the same techniques and insights, and are best described using the example of a CPU-initiated RPC using addresses homed on the device (4(a)).
+The protocols use the same techniques and insights, and are best described using the example of a CPU-initiated invocation on the accelerator using addresses homed on the device (5(a)).
 An implementation of this protocol is used for the evaluation in Section 6.1.
+Note that, unlike RPC, invocations are local to the machine and avoid marshaling or unmarshaling of messages.
   1. (1)
-    The protocol starts with line B in Exclusive state in the CPU cache, and line A Invalid.
+    The protocol starts with line B in Exclusive state in the CPU cache, and line A Invalid (we show how this can be conveniently achieved below).
 To initiate an operation, the CPU writes the request data from its registers into line B, which transitions to Modified.
@@
   5. (5)
-    The CPU cache receives line A’s new data and the CPU unblocks, immediately loading the first word of the RPC response into a register.
-As described, the resulting state is that A is now Invalid and B is Shared in the CPU cache.
+    The CPU cache receives line A’s new data and the CPU unblocks, immediately loading the first word of the invocation response into a register.
+As described, the end state is that A is now Invalid and B is Shared in the CPU cache.
 However, in some deployed MESI-like protocols (including the Enzian interconnect we use), the device is free to return the data stipulating that it is not Shared but now Exclusive to the CPU.
 This final feature leaves the cache in the state it started in, with the roles of A and B reversed, ready for the next transaction.
-The whole exchange takes 2 round trips on the coherent interconnect.
+The whole exchange takes 2 round trips on the coherent interconnect, and require no spinning on the CPU.
 The other 3 protocols work similarly.
-For example, 4(b) shows the analogous case where the device initiates the request to the CPU, but still using lines homed on the device.
+For example, 5(b) shows the analogous case where the device initiates the request to the CPU, but still using lines homed on the device.
 This is used for the NIC experiment in Section 6.3.
@@
 In this paper we therefore focus on the case where the cache lines are homed on the device.
-4.4. Discussion
-A number of concerns might be raised about these protocols in practice; we address them in this section.
-First is the issue of timeouts: we are blocking a core’s request for a cache line until a computation has completed.
-If this triggers a timeout in the CPU’s cache, the result is likely to be a memory fault exception, a machine check, or the processor simply locking up (on the Enzian, it is a machine check).
-We can work around these issues in practice by a small state machine on the device return a “not ready yet” response in the event of an approaching timeout, causing the software on the issue another request for the other cache line, extending the response time effectively indefinitely without the need for spinning.
-In practice, timeouts are quite long and the computations we target are short.
-The second issue is more subtle, and concerns possible deadlock due to the device stalling a request until it can issue, and gets a reply back from, another to a different cache line.
-Such standards as exist for interoperable cache coherent interconnects are silent on what happens in this situation, possibly because it has not occurred until now.
-The design assumes that, in the platform, transactions on different cache lines can progress independently.
+Subtleties:
+A number of issues must be handled to make this practical, starting with timeouts: we block a core’s request for a cache line until a device operation has completed.
+A timeout in the CPU’s cache is likely to cause a memory fault exception, a fatal machine check, or the processor simply locking up (on Enzian, it is a machine check).
+We solve this issue with a small state machine on the device that returns a “not ready yet” response before an impending timeout, causing the software on the issue another request for the other cache line, extending the response time indefinitely without the need for spinning.
+In practice, we target operations that are shorter than the timeout.
+The second issue concerns deadlock due to the device stalling a request until it can issue, and get a reply back for, another request to a different cache line.
+Existing standards for interoperable cache coherent interconnects are silent on what happens in this situation, possibly because it has not occurred until now.
+The design assumes that transactions on different cache lines can progress independently.
 In most implementations, this assumption typically holds in order to maximize the memory bandwidth utilization and minimize request latency, and also to simplify reasoning about deadlock freedom of the coherence protocol itself.
@@
 Were this to be the case, it could be avoided by careful placement of A and B in the physical address space.
-This is actually the case in our FPGA device implementation, described in Section 4.5.
-4.5. Implementation in Enzian
-We implemented both device-homed protocols in the Enzian platform, which we use for the evaluation in Section 5.
-The implementation uses SystemVerilog and exploits an existing “Directory Controller” on Enzian’s FPGA (Ramdas, 2024) which manages the coherent links to the CPU and handles transient states in the ECI protocol.
+This is actually the case in our FPGA device implementation, described in Section 5.
+Implementation in Enzian
+We implemented the two device-homed protocols over Enzian and evaluate them in Section 6.
+The implementations use SystemVerilog and SpinalHDL (contributors, 2024) and exploit an existing “Directory Controller” on Enzian’s FPGA (Ramdas, 2023) which manages the coherent links to the CPU and handles transient states in the ECI protocol.
 ECI provides all the messages we require to implement our messaging protocols.
 In particular, it includes the ability for the device to return a line to the CPU cache in Exclusive state, even though the CPU requested it in Shared.
-This optimization (discussed in Section 4.3) dramatically reduces latency, as shown in Section 5.1.
+This optimization dramatically reduces latency, as shown in Section 6.1.
 In our case, no flow-control issues arise since ECI uses separate virtual channels for request and reply messages, removing the chance of deadlock.
@@
 A GSync blocks any subsequent transactions until all those started before the GSync have retired.
-But the invalidates sent are not awaited by the core they are still completed allowing the protocol to progress.
+Fortunately, any invalidations sent are not blocked by core, allowing the protocol to progress regardless.
 The ThunderX-1 divides its last-level cache functionality into a set of units called TADs, each of which can handle up to 16 simultaneous cache transactions.
-Two consecutive cache lines will be mapped to the different TADs by L2 to ensure that the operations can proceed independently.
-On this CPU, barriers are needed to reliably order reads and writes.
-For example, it is critical that the core’s write buffer is drained into L1 cache (which is write-through to the L2) before a subsequent load from the core signals to the FPGA that the line can be pulled.
-The ARM architecture’s weak memory model requires a DMB barrier instruction between the writes and the read, and on the ThunderX-1 this is sufficient to ensure ordering observed at the FPGA.
-On the FPGA, control (interception) is performed based on AXI read requests, since both Shared and Exclusive operations trigger these requests.
-The AXI writes go directly to the application.
-If the read is the result of the Exclusive, then the dummy data is sent back to the CPU as here is no fresh result yet.
+We ensure that consecutive cache lines are mapped to different TADs to ensure that the operations can proceed independently.
+The ARMv8-A weak memory model means that barriers are needed to reliably order reads and writes.
+For example, it is critical that the core’s write buffer is drained into L1 cache (which is write-through to the L2) before a subsequent load from the core signals to the FPGA that the line can be pulled, and this requires a DMB barrier instruction between the writes and the read.
+On the ThunderX-1 this is sufficient to ensure ordering observed at the FPGA.
 6. Evaluation
-In this section, we compare PIO over a coherent interconnect with two other reference points: PIO directly over PCIe– essentially, the CPU copying data to and from a region of PCIe address space – and descriptor-based DMA over PCIe.
+We compare PIO over a coherent interconnect with two other reference points: PIO directly over PCIe– essentially, the CPU copying data to and from a region of PCIe address space – and descriptor-based DMA over PCIe.
 All use the same Enzian hardware platform.
@@
 DMA experiments all use the Xilinx XDMA hardware IP on the FPGA.
-We explore three different application scenarios: synchronous RPC to an accelerator on the FPGA, a closely-coupled high-speed NIC, and a stream processor which moves some of its dataflow operator graph to the FPGA.
-5.1. RPC to an accelerator
-The first experiment measures null-RPC latency when the CPU invokes a function on the device, passing arguments, and obtaining a result.
-The protocol shown in 4(a) and detailed in Section 4.3 is used (i.e.
-the cache lines used for communication are homed on the FPGA).
-Figure 5. Round trip ECI RPC latency, plus software latency between two sockets
+We explore three different application scenarios: synchronous invocation to an accelerator on the FPGA, a closely-coupled high-speed NIC, and a stream processor which moves some of its dataflow operator graph to the FPGA.
+6.1. Invoking functions on an accelerator
+We first measure latency when the CPU invokes a function on the device, passing arguments, and obtaining a result. We use the protocol shown in 5(a), whereby the cache lines are homed on the FPGA.
+Figure 6. Invocation latency over ECI vs. FastForward
 Figure 6 shows the distribution of total latency for the this operation.
@@
 When we introduce the optimization of returning the line in Exclusive, the median latency drops to around 900ns.
-For comparison, we also show an implementation of the FastForward protocol (Giacomoni et al., 2008a) exchanging cache lines between sockets in a dual-socket ThunderX-1-based Gigabyte R150-T61 server with similar CPU and DDR specification to Enzian; this achieves a median latency of about 1150ns.
-The optimized protocol is performing 2 round-trip message exchanges over ECI, which has a one-way latency at the link later of about 150ns.
+For comparison, we also show an implementation of the FastForward protocol (Giacomoni et al., 2008) exchanging cache lines between CPU sockets in a dual-socket ThunderX-1-based Gigabyte R150-T61 server with similar CPU and DDR specification to Enzian; this achieves a median latency of about 1150ns.
+The optimized protocol is performing 2 round-trip message exchanges over ECI, which has a one-way latency at the link layer of about 150ns.
 The rest of the overhead (300ns) is mostly incurred in the protocol processing in the directory controller.
 The FPGA is clocked at about 300MHz.
-A significant penalty is incurred when the results of the RPC are returned in Shared, resulting in further round trips over ECI to (unnecessarily) invalidate the line at the FPGA before sending the next message.
-This suggests that future coherent interconnects would benefit greatly from allowing this kind of operation.
+A significant penalty is incurred when the results of the call are returned in Shared, resulting in further round trips over ECI to (unnecessarily) invalidate the line at the FPGA before sending the next message.
+This suggests that future coherent interconnects would benefit greatly from allowing the home node to return a line in Exclusive.
 Overall, the optimized protocol comfortably outperforms FastForward, despite the latter exploiting highly optimized hardware implementations at both ends (clocked at 2.1GHz) and aggressively polling, whereas our optimized protocol requires no spinning on the part of the CPU or device.
-5.2. RPC over ECI vs. PCIe PIO and DMA
-We now compare the RPC latency achieved using our protocol over ECI with PIO over PCIe and the Xilinx XDMA hardware using descriptors.
-{listing}
-[t] {minted} [ frame=lines, framesep=2mm, baselinestretch=1.2, bgcolor=LightGray, fontsize=, linenos, ] c // map 4KiB of PCIe memory for read/write int fd = open(”/sys/bus/pci/devices/…/resource0”, O_RDWR); void *pcie_bar = mmap(NULL, 4096, PROT_READ — PROT_WRITE, MAP_SHARED, fd, 0);
-// write RPC arguments to PCIe memory *(volatile rpc_args_t *)pcie_bar = rpc_args; __sync_synchronize();
-// read RPC result from PCIe memory rpc_res_t rpc_res = *(volatile rpc_res_t *)pcie_bar; __sync_synchronize(); Example RPC using PIO over PCIe on Linux. Error handling not shown for brevity.
-For PIO over PCIe we pre-map the relevant PCIe aperture (BAR) into user space on the CPU and measure the time to write a value to this space and read a result back, as illustrated in Section 5.2.
-For XDMA numbers we use Xilinx XDMA benchmark (Xilinx DMA Benchmark, 2024), which essentially performs the same operation using descriptor-based DMA.
-In both these cases, the FPGA is configured to map the request and response into a write to and a read from onboard Block RAM, respectively.
+6.2. Invocation over ECI vs. PCIe PIO and DMA
+Figure 7. Invocation latency for different payload sizes
+We now compare the latency achieved in the previous section with both PIO over PCIe and the Xilinx XDMA hardware using descriptors.
+For DMA numbers we use the Xilinx XDMA benchmark (Xilinx DMA Benchmark, 2024), which essentially performs the same operation using descriptor-based DMA.
+In both these cases, the FPGA is configured to map the request and response into a write to, and a read from, onboard Block RAM.
 As in Section 6.1, we measure latency to send a request from the CPU to the FPGA and read a response.
@@
 The results are shown in Figure 7.
-Up to payloads of 8KiB, the latency of XMDA-based communication is completely dominated by descriptor setup and manipulation, and remains largely independent of payload size.
-Faster PCIe interconnects *recall that this is a PCIe Gen3 x8 link) would only further emphasize this.
+Up to payloads of 8KiB, the latency of XMDA-based communication is dominated by descriptor setup and manipulation, and remains largely independent of payload size.
+Faster PCIe interconnects (this is a PCIe Gen3 x8 link) would amplify this effect.
 Latency over ECI is dramatically lower for smaller payloads.
 Unsurprisingly, it is constant up to 128 bytes (the size of a cache line on Enzian) and then increases linearly, since the protocol simply replicates each cache line exchange until all the data has been transferred.
-It is only at 8KiB that ECI RPC latency exceeds that of XMDA.
-One might argue that the transfer of multiple cache lines over ECI might be improved by clever pipelining and locating the lines in physical memory so as to exploit parallelism in the directory controllers, but we do not explore that in this paper.
-Figure 6. RPC latency for different payload sizes
-[…]
-On this platform, using common techniques to improve performance like vector instructions has limited or no benefit for PIO over PCIe.
-For writes, the hardware already performs coalescing, and we see transactions of up to 512 bits regardless of vector size (confirmed using a logic analyzer).
-For reads, the 128 bit limit remains, and is responsible for the increase in latency between 16-byte and 32-byte payloads.
+It is only at 8KiB that ECI invocation latency exceeds that of XMDA.
+Arguably the transfer of multiple cache lines over ECI might be improved by clever pipelining and locating the lines in physical memory so as to exploit parallelism in the directory controllers, but we do not explore that in this paper.
+Relative to ECI, PIO over PCIe performs poorly except for payloads under 16 bytes in size. This is due to limitations in PCIe that we have discussed in Section 4, notably the limited read bus width.
 We conclude that up to the page size (4KiB) or the typical size of jumbo frames, PIO remains significantly faster than the DMA option commonly used for these transaction sizes.
 Thus even using the XDMA implementation for the experiments we still have the room for comparison with optimized DMA implementations.
-5.3. Network interface
-[…]
-We implement 3 variations of a 100Gb/s NIC on Enzian.
-The first (7(a)) is a conventional approach connecting one of Enzian’s 100Gb Ethernet MACs fairly directly to the XDMA engines, resulting in a NIC which is programmed using descriptor rings.
-The second design (7(b)) buffers packets in onboard SRAM on the FPGA and allows the CPU to read and write packet and control data using PCIe reads and writes.
-The final approach (7(c)) uses a custom module on the FPGA to bridge between the ECI directory controller and the Ethernet MAC, using the protocol in 4(b) to deliver network packets to the CPU in cache lines.
-In all cases we clock the core NIC logic at 250 MHz.
+This experiment is closed-loop and single-threaded, and so achieved throughput is proportional to the round-trip latency.
+Fully saturating the raw link bandwidth (in all cases) would require many parallel invocations at once, which would go against our goal of supporting applications where latency is paramount.
+6.3. Network interface adapter
 (a) PCIe XDMA: packets transferred using descriptor-based DMA
@@
 Figure 8. Different NIC architectures. Blue denotes data movers; red denotes memory that holds packet data.
+[…]
+We implement 3 variations of a 100 Gb/s NIC in Enzian.
+The first (8(a)) is a conventional approach connecting one of Enzian’s 100 Gb/s Ethernet MACs directly to the XDMA engines, resulting in a NIC programmed using descriptor rings.
+The second design (8(b)) buffers packets in onboard SRAM on the FPGA and allows the CPU to read and write packet and control data using PCIe reads and writes.
+The final approach (8(c)) uses a custom module on the FPGA to bridge between the ECI directory controller and the Ethernet MAC, using a slight variant of the protocol in 5(b) to deliver network packets to the CPU in multiple cache lines.
+In all cases we clock the core NIC logic at 250 MHz.
+Figure 9. NIC implementations: receive latency
+Figure 10. NIC implementations: transmit latency
+Table 1. NIC implementations: latency percentiles on selected packet sizes. Higher tail latencies are marked in bold.
+Size Percentiles RX ($\mu$s) Percentiles TX ($\mu$s)
+P50 P95 P99 P100 P50 P95 P99 P100
+PCIe DMA
+64 65.39 66.36 67.65 100.01 10.06 10.35 10.59 16.49
+1536 64.77 65.59 66.21 133.84 10.89 11.12 11.33 30.84
+9600 65.89 67.12 68.05 123.61 15.73 16.05 16.29 41.99
+PCIe PIO
+64 3.25 3.26 3.27 3.39 0.34 0.36 0.38 4.80
+1536 72.89 72.93 72.96 73.05 1.82 1.84 1.86 6.40
+9600 450.28 450.35 450.38 451.10 9.91 10.01 10.07 10.14
+ECI PIO
+64 1.05 1.06 1.07 1.17 1.06 1.12 1.14 1.18
+1536 7.24 7.29 7.39 7.43 3.09 3.26 3.50 3.59
+9600 39.43 39.48 39.50 39.55 9.07 9.19 9.65 9.95
 For experiments we configure the Ethernet MAC in near-end PCS/PMA loopback mode for a reliable measurement of packet delivery inside the server without network delays.
@@
 We do not include the time it takes the packet to go through the Ethernet PCS/PMA loopback, since the MAC is fixed and the same in all implementations presented in this experiment.
+To minimize disruptions from the Linux scheduler, we follow the common practice to free one CPU core from kernel tasks and IRQ processing (Akkan et al., 2012), with the isolcpus function together with core pinning using taskset.
+In this setup, the only disturbance from Linux would be the 250 Hz timer interrupt present on all cores.
+While a custom, completely tickless kernel would remove this disturbance completely, it is not commonly deployed in production environments due to performance complications.
+We use such a tickless kernel to measure tail latency below, but other experiments use a stock generic kernel from Ubuntu.
 We show figures for XDMA in both polled and interrupt-driven modes; in practice the difference in latency is small.
 Figure 9 shows receive latency for different packet sizes for each of the NIC implementations.
-As in the RPC experiment, DMA latency is fairly stable across all packet sizes, suggesting that the latency is dominated by various DMA overheads, such as descriptor setup and manipulation.
-Note that system calls here actually result in very little overhead (only a few microseconds); the overhead is actually dominated by cache misses manipulating DMA data structures in main memory.
-[…]
+As in the invocation experiment, DMA latency is fairly stable across all packet sizes, suggesting that the latency is dominated by various DMA overheads, such as descriptor setup and manipulation.
+Note that system calls here actually result in little overhead (only a few $\mu$s); the overhead is dominated by cache misses manipulating DMA data structures in main memory.
+In contrast, PIO over both ECI and PCIe offer much lower latency for small packet sizes up to 1024 bytes DMA, highlighting high efficiency gains in PIO compared to DMA.
+However, the latency of PCIe PIO quickly degrades for larger packets due to the non-posted read problem discussed in Section 4, reaching over 350 $\mu$s for 8 KiB packets.
+In contrast, ECI offers dramatically better latency even at 8 KiB packets.
 Figure 10 shows transmit latency.
-In this case, PCIe PIO competes well with ECI across all packet sizes due to combining posted writes.
+In this case, PCIe PIO competes well with ECI across all packet sizes due to combining posted writes as we discussed in Section 4.
 PIO over ECI shows a slightly higher latency for small packets, due to it needing 2 round-trips over ECI for each 128 bytes, versus a single PCIe round trip.
 Both PIO solutions perform significantly better than DMA across all packet sizes.
-These results show a clear advantage of using PIO over DMA, even for larger packets.
-The poor receive performance of PCIe-based PIO for receiving packets is largely due to a limitation of the PCIe design, that disappears for transmit.
-5.4. Worked Example: Offloading Timely Dataflow
-In this example we apply the techniques and results outlined so far to a practical workload: Timely Dataflow (Timely Dataflow, 2020), a real-time stream-processing system.
+Table 1 shows the tail latency impact for selected packet sizes for each of the NIC implementations.
+We select three representative packet sizes to evaluate tail latency: 64 for the smallest possible packet, 1536 for the normal Ethernet MTU, and 9600 for a common jumbo frame MTU.
+We use the tickless kernel setup mentioned earlier to avoid the 250 Hz Linux timer, in order to reveal the actual tail latency impact from our NIC.
+We see that PIO over PCIe results in low tail latency compared to the descriptor-based approach, while ECI completely eliminates tail latency in this scenario.
+These results show a clear advantage of using PIO over DMA, even for larger packets, and especially when using a coherent interconnect .
+The poor receive performance of PCIe-based PIO for receiving packets is due to the unposted-read limitation of PCIe, and disappears for transmit.
+6.4. Accelerating Timely Dataflow
+In this example we apply the techniques and results outlined so far to a practical workload: Timely Dataflow (Murray et al., 2013), a real-time stream-processing system.
 We compare the baseline performance to a partially FPGA-offloaded implementation under three CPU–FPGA communication scenarios: DMA over PCIe, PIO over PCIe, and PIO over ECI.
@@
 Operators may be arbitrarily complex and are independent, and thus benefit from offloading where the overhead of shipping the data to the FPGA is compensated by greater throughput of the offloaded operators.
-Figure 8. NIC implementations: receive latency
-Figure 9. NIC implementations: transmit latency
 To test the limits, we construct a worst-case scenario: a large (31-element) graph of simple operators (filters).
 The large graph maximizes communication overhead in the form of progress tracking data, while the simple operators leave the FPGA with only a small performance lead over the CPU.
-We want to predict whether PIO can make FPGA offload work in regimes where the fixed overheads of DMA make it impractical, for example where small batch sizes are needed to ensure data freshness (as in Differential Dataflow).
+We want to predict whether PIO can make FPGA offload work in regimes where the fixed overheads of DMA make it impractical, for example where small batch sizes are needed to ensure data freshness (as in Differential Dataflow (McSherry et al., 2013)).
 The first (blue) bars of Figure 11 show the baseline performance of the non-offloaded (CPU-only) system against batch size.
@@
 As shown by the $r^{2}$ values, the fit for each dataset is excellent.
-We can thus model the time taken to process a batch as a fixed component, or overhead, independent of the amount of data processed (e.g.
-running the Timely scheduler, or setting up DMA descriptors) plus a component which scales linearly with the size of the batch.
+We can thus model the time taken to process a batch as a fixed component, or overhead, independent of the amount of data processed (e.g., running the Timely scheduler, or setting up DMA descriptors) plus a component which scales linearly with the size of the batch.
 Compared to the baseline, DMA offload has a significantly lower per-byte cost ($3.7$ vs.
@@
 As the hardware can handle many in-flight transfers at once, modifying the protocol to transfer 32 lines at once would reduce the per-byte cost enough that ECI-PIO offload breaks even.
-Alternatively or additionally, bringing the fixed cost down reduces the smallest practical batch.
-If the relative overhead could be brought below 500ns/batch then with parallel cache lines giving a 1ns/B transfer cost, FPGA offload using ECI-PIO would outperform the CPU from a batch size of 128B, or a single cache line.
-With such an optimized protocol, ECI-PIO would outperform all other cases until a batch size of 142KiB, above which PCIe-DMA would overtake it giving a net offload benefit across all batch sizes.
-Figure 10. DMA and PIO performance for Timely Dataflow FPGA offload with a graph size 31
+Additionally, bringing the fixed cost down reduces the smallest practical batch.
+If the relative overhead could be brought below 500ns/batch then with parallel cache lines giving a 1ns/B transfer cost, FPGA offload using ECI-PIO would outperform the CPU from a batch size of 128B.
+With such an optimized protocol, ECI-PIO would outperform all other cases until a batch size of 142KiB, above which PCIe-DMA would overtake it.
+Figure 11. Timely Dataflow offload, graph size 31
+Table 2. 1st-order (offset + slope) coefficients fitted to the data series of Figure 11, as cost per byte and fixed overhead
 Time/Byte (ns) Overhead ($\mu$s) $r^{2}$
@@
 ECI-PIO 30 221 1.00
-Table 1. 1st-order (offset + slope) coefficients fitted to the data series of Figure 10, as cost per byte and fixed overhead
+6.5. Offloading Bloom Filters
+In a second acceleration experiment we offloaded Bloom filters (Bloom, 1970) to the FPGA.
+Bloom filters are used to efficiently check whether an element is a member of a set.
+It uses little memory and allows for fast lookups.
+A Bloom filter is equipped with $k$ different hash functions and a lookup table.
+Testing the presence of an element requires computing these $k$ hash functions and querying the lookup table with the results.
+We took inspiration from the applications discussed in Fleet (Thomas et al., 2020).
+As the CPU baseline, we re-implemented the Fleet C++ code in Rust as a Timely Dataflow operator.
+In the Fleet implementation, the hashes were computed in parallel for each byte sequentially using AVX instructions, which we replicated in Rust using SIMD instructions.
+On the FPGA side, the hash computation is parallelized and the iteration over the element’s bytes is pipelined.
+We present the numbers of using a single thread on CPU and a single compute unit on FPGA in order to keep the comparison fair.
+With the Bloom filter integrated into Timely Dataflow as a single offloaded operator, most of the additional overhead from processing sub-graphs is eliminated and the flow of data between FPGA and CPU is reduced to the minimal practical batch.
+In addition, the computational load is significantly higher compared to a simple filter.
+The $k$ hash functions consist of sequential bit shifting, addition, and XOR-ing, which are very fast in FPGA logic and deliver the offloading benefit comes from.
+Therefore in this application we concentrate on minimal practical batches and direct offloading benefits.
+We implement the Bloom filter on the FPGA to take 128-byte elements with $k=8$ hash functions that take byte-wide inputs.
+We unroll the hash function calculations for each byte lane by a factor of 2, resulting in a 64-cycle latency for each element.
+We then pipeline the calculations with an initiation interval of 2 cycles to saturate the 512-bit bus.
+The return value is the set of 8 64-bit hashes for each element.
+Instead of directly returning the hash results, the FPGA could have computed the lookup table positions or performed the lookup itself, returning either $k$ indices or binary values.
+In the case of indices, the read-to-write ratio would stay the same as in our approach, but for the lookup, the read requests would be minimized and writes would dominate.
+In this experiment, we assume that the lookup table is located elsewhere and focus on efficiently calculating the hash function and transferring the results to the lookup table.
+In Figure 12, the latencies for PIO and CPU both start around 25 $\mu$s, due to the high overhead of streaming the input data.
+The real processing time of one element is 2.6 $\mu$s on the CPU and 1.7 $\mu$s when offloading with ECI.
+As overall processing time increases with the data size, we see the gap between CPU and PIO more clearly.
+After 8 KiB, same as in previous experiments, XDMA starts to perform better than other approaches as performance becomes dominated by the higher bulk throughput of PCIe DMA.
+Figure 12. Offloading Bloom filters to the FPGA
 7. Related work
+Cho et al.
+(Cho et al., 2018) provide an excellent survey of existing CPU-device communication mechanisms, and address high-speed storage devices with microsecond latencies.
+They question the conventional wisdom that hiding the latency of interacting with such devices is not possible using existing techniques, and show how access latency using existing descriptor-based protocols can be effectively hidden by clever use of prefetching, better hardware queues, and user-level context switching.
+Our work is different (and complementary) in reducing the actual latency of each message between CPU registers and device memory.
+7.1. Improving descriptor-based DMA
+Cohort (Wei et al., 2023b) proposes a single, uniform, single-producer, single-consumer, queue-based DMA descriptor interface to all accelerators on an System-on-Chip (SoC) which replaces the variety of ad-hoc interfaces found on most SoCs.
+Cohort mandates an MMU on each accelerator, which maintains consistency with CPU MMUs and allows seamless user-space access to accelerator queues, and integrates with the cache coherence protocol to optimize the use of lock-free descriptor queues.
+We focus instead on what a latency-optimized interface might look like, eschewing descriptors in favor of direct cache line transfers.
+Ensō (Sadok et al., 2023) shows the benefits of replacing the packet-based DMA software interface used by modern NICs with one based on a contiguous large buffer of opaque, variable-sized messages, with metadata sent via a different completion buffer.
+Ensō replaces descriptors with this buffer using an ingenious combination of direct MMIO writes to the PCIe device and polling memory queues written by the NIC.
+Our work similarly tries to reduce the overhead of descriptor management, but by focusing solely on latency-sensitive small payloads and avoiding DMA access completely.
 7.2. Coherent interconnects
@@
 It and the competing Gen-Z (Keeton, 2015) and CCIX (CCIX Consortium and others, 2024) protocols have meanwhile been either merged into or replaced by CXL (Li et al., 2023), which seems likely to be the standard interoperable protocol in the short- to medium-term future.
-Commercial CXL hardware implementing the more recent revisions, which allow less restricted use cases, is still not widely available.
-Consequently research on the software uses of coherent interconnects is still largely performed on specifically research-oriented machines such as the Intel HARP (Intel, 2024; Intel Harp, 2024), and Enzian (Cock et al., 2022).
-While HARP provides a pre-configured coherent cache in the FPGA with a conventional interface, Enzian’s ECI protocol is fully open and customizable.
+However, commercial CXL hardware implementing the more recent revisions, which allow less restricted use cases, is still not widely available.
 Other notable coherent interconnects include the RISC-V-specific TileLink (Berkeley Architecture Research, 2022; Terpstra, 2017; Cook et al., 2017) and NVIDIA’s NVLink 2.0 (NVIDIA, 2024; Wei et al., 2023a).
@@
 Existing work has built on the available research platforms to explore the design space for practical applications.
-Centaur (Owaida et al., 2017) demonstrated the FPGA offload of database operations on Intel HARP v1, and Cohort (Wei et al., 2023b; Mukherjee et al., 1996) proposed and prototyped a unified queue-based software interface to such accelerators building on the P-Mesh interconnect of OpenPiton (OpenPiton Open Source Research Processor, 2024).
+Centaur (Owaida et al., 2017) demonstrated the FPGA offload of database operations on Intel HARP v1.
+However, while Intel HARP (Intel, 2024; Intel Harp, 2024) coupled a CPU and FPGA using a coherent QPI link, it provided a pre-configured coherent cache in the FPGA rather than user access to coherence protocol messages.
+In contrast, this access is readily available in Enzian (Cock et al., 2022).
 Open, extensible protocols have been used to explore coherent offload in simulation, including making the case for protocol specialization building on the Spandex protocol family (Alsop et al., 2022, 2018).
@@
 7.3. Cache line-based communication
-Other work (Tamimi et al., 2022) has already noted that the cache line is a better unit of transfer where small operations are expected.
-This is a common observation, long since applied in the original FastForward (Giacomoni et al., 2008b) and Barrelfish UMP (Baumann et al., 2009b) protocols.
+Communication in software between peer modes in a cache-coherent system is typically very different from using descriptor-based DMA.
+FastForward (Giacomoni et al., 2008), Barrelfish (Baumann et al., 2009), Concord (Iyer et al., 2023), and Shinjuku (Kaffes et al., 2019) adopt a much more direct approach to sending small messages with low latency, e.g., by exploiting the cache coherence protocol to transfer lines between caches on demand, and local polling to provide synchronization.
+Recently, cache coherence between NUMA CPUs has been used to simulate communicating between software and a hypothetical cache-coherent NIC (Schuh et al., 2024).
+Other work (Tamimi et al., 2022) has already noted that the cache line is a better unit of transfer where small operations are expected, as in FastForward (Giacomoni et al., 2008) and Barrelfish (Baumann et al., 2009) protocols.
 More recently Concord (Iyer et al., 2023) communicates scheduling decisions between workers and the dispatcher via a polled cache line, converting worker threads from interrupt-driven to poll-mode “CPU drivers”.
@@
 A recent study of data center RPC from Google (Karandikar et al., 2021) reinforces the importance of small transfers and highlights the large latency of PCIe transactions.
+HyperPlane (Mirhosseini et al., 2020) observes that I/O stacks in cloud data centers frequently resort to spin polling the many descriptor queues provided by modern NIC hardware, and proposes a a new QWAIT hardware instruction which leverages the processor’s cache mechanism (much like MWAIT) to watch many different in-memory descriptor queues at once, providing hardware acceleration for select() or epoll()-like operations.
+Evaluation is performed using the Gem5 simulator, although this implementation is not available.
+We take a different approach based on transferring data directly in cache lines without descriptors.
 7.4. Exploring PIO
-The space where PIO is preferable to DMA is likewise well studied.
+Neugebauer et.
+al.
+(Neugebauer et al., 2018) provide a comprehensive analysis of PCIe performance in the context of NICs.
+The space where PIO is preferable to DMA for PCIe has been explored before.
 The hXDP (Brunella et al., 2020) FPGA NIC work highlighted the high overhead of small PCIe transactions, and consequently performs small batch computations solely on the CPU.
@@
 Compared to a traditional DMA NIC, their FPGA-based prototype shows better latency and throughput for small and medium-sized messages and comparable throughput for large messages.
-The ThunderX CPU automatically performs combining for PCIe writes, and the results of Section 5.2 thus reflect this optimization.
-Neugebauer et. al. (Neugebauer et al., 2018) provide a comprehensive analysis of PCIe performance in the context of NICs.
+The ThunderX-1 automatically performs combining for PCIe writes, and the results of Section 6.2 thus reflect this optimization.
 Dagger (Lazarev et al., 2021) builds on CCI-P, the commercialized implementation of HARP’s coherent cache, to construct an FPGA-based NIC specialized for low latency RPC.
 A host-coherent cache holds connection states and the necessary structures for the transport layer on the NIC, while the payload remains in host memory.
-This minimizes FPGA memory demands, and exploits the lower latency of cache misses compared to conventional PCIe-based NICs.
+This minimizes FPGA memory demands, and exploits the lower latency of cache misses compared to conventional PCIe-based NICs..
 8. Conclusion and Future Work

### v3

@@
 Abstract.
-Conventional wisdom holds that an efficient interface between an OS running on a CPU and a high-bandwidth I/O device should be based on Direct Memory Access (DMA), descriptor rings, and interrupts: DMA offloads transfers from the CPU, descriptor rings provide buffering and queuing, and interrupts facilitate asynchronous interaction between cores and device with a lightweight notification mechanism.
-In this paper we question this wisdom in the light of modern hardware and workloads, particularly in cloud servers.
-Like others before us, we argue that the assumptions that led to this model are obsolete, and in many use-cases use of Programmed I/O (PIO), where the CPU explicitly transfers data and control information to and from a device via loads and stores, actually results in a more efficient system.
-However, unlike others to date, we push this idea further and show, in a real implementation, the gains in average and tail latency for fine-grained communication achievable using an open cache-coherence protocol which exposes cache transitions to a smart device.
-We show this using three use-cases: fine-grained RPC-style invocation of functions on an accelerator, offloading of operators in a streaming dataflow engine, and a network interface targeting for serverless functions, comparing our use of coherence with both traditional DMA-style interaction and a highly-optimized implementation using PIO over PCI Express (PCIe).
+Conventional wisdom holds that an efficient interface between an OS running on a CPU and a high-bandwidth I/O device should use Direct Memory Access (DMA) to offload data transfer, descriptor rings for buffering and queuing, and interrupts for asynchrony between cores and device.
+In this paper we question this wisdom in the light of two trends: modern and emerging cache-coherent interconnects like CXL3.0, and workloads, particularly microservices and serverless computing.
+Like some others before us, we argue that the assumptions of the DMA-based model are obsolete, and in many use-cases programmed I/O, where the CPU explicitly transfers data and control information to and from a device via loads and stores, delivers a more efficient system.
+However, we push this idea much further.
+We show, in a real hardware implementation, the gains in latency for fine-grained communication achievable using an open cache-coherence protocol which exposes cache transitions to a smart device, and that throughput is competitive with DMA over modern interconnects.
+We also demonstrate three use-cases: fine-grained RPC-style invocation of functions on an accelerator, offloading of operators in a streaming dataflow engine, and a network interface targeting serverless functions, comparing our use of coherence with both traditional DMA-style interaction and a highly-optimized implementation using memory-mapped programmed I/O over PCIe.
 1. Introduction
-Modern interfaces between CPUs and high-performance devices like network interface adaptors (NICs), storage controllers, Graphics Processing Units (GPUs), and other computational accelerators are designed to optimize throughput.
-Built over a PCIe interconnect, descriptor rings in main memory hold queues of operations and completions, and the device performs DMA transactions to read and write both descriptors and payload data.
+Modern interfaces between CPUs and high-performance devices like network interface adaptors (NICs), Graphics Processing Units (GPUs), etc.
+are designed to optimize throughput for large transfers.
+Built over a PCI Express (PCIe) interconnect, descriptor rings in memory hold queues of requests and completions, and the device uses DMA to read and write both descriptors and payload data.
 This design, along with the underlying PCIe interconnect, trades off latency for small transactions in favor of high throughput for large ones.
-In this paper we question this consensus in the light of workloads which depend for performance not on bulk throughput, but the cumulative latency of small transactions between CPU and device: closely-coupled hardware accelerators for irregular workloads, or data center Remote Procedure Call (RPC) using small messages.
-We do this in the context of emerging cache-coherent peripheral interconnects like CXL.cache 3.0 (CXL Consortium, 2024).
-The continuing increase in PCIe bandwidth means that, for a given transfer size, the significance of PCIe latency (time to first byte) has become more of a concern, leading to many proposals for reducing the overhead of descriptor management (which we survey in Section 7.
-However, we propose a more radical approach, aimed at use-cases where reduced latency is more important than higher maximum throughput: using PIO to Memory-Mapped I/O (MMIO) devices directly, and avoiding DMA, descriptor queues, operation buffering, and interrupts entirely.
-We use a real hardware platform which implements a coherent interconnect between a server-class CPU and a large FPGA to investigate the trade-offs between DMA using descriptor rings, PIO directly over a PCIe interconnect, and PIO over a full cache-coherence protocol.
-We show that, for small (around 1KiB or less) transfers, PIO posted writes over PCIe significantly outperform DMA, although reads are less efficient since, unlike writes, they cannot be pipelined and so incur PCIe’s significant round-trip latency penalty.
-However, a full cache coherence protocol which avoids PCIe significantly outperforms both DMA and PIO over PCIe.
-Crucially, not only is median latency considerably lower, but the tail of the latency curve is reduced significantly when using PIO over PCIe, and completely eliminated using the cache coherence protocol.
-We go further, however, and show how a conventional MESI-like coherence protocol can be exploited in novel ways by intelligent devices which can interact with an unmodified CPU at the coherence message level.
+We question this consensus in the light of workloads which depend for performance not on bulk throughput, but the cumulative latency of small transactions between CPU and device: closely-coupled accelerators for irregular workloads, or Remote Procedure Call (RPC) using small messages.
+We do this in the context of emerging cache-coherent peripheral interconnects like CXL.mem 3.0 (CXL Consortium, 2024).
+The continuing rise in PCIe bandwidth means, for a given transfer size, the significance of PCIe latency (time to first byte) is now more of a concern, leading to many proposals for reducing the overhead of descriptor management (which we survey in Section 6).
+We propose a more radical approach, aimed at use-cases where low latency is more important than high maximum throughput: using Programmed I/O (PIO) to issue loads and stores to Memory-Mapped I/O (MMIO) registers directly, avoiding DMA, descriptors, queues, and interrupts entirely.
+Using a real hardware platform which implements a coherent interconnect between a server-class CPU and a large FPGA, we investigate the trade-offs between DMA with descriptors, PIO directly over a PCIe interconnect, and PIO over a full cache-coherence protocol.
+We show that, for small ($\leq$ 1KiB) transfers, PIO posted writes over PCIe significantly outperform DMA, although reads are much less efficient since, unlike writes, they cannot be pipelined and so incur PCIe’s significant round-trip latency, in line with other recent critiques of DMA (Liew et al., 2025; Sadok et al., 2023; Schuh et al., 2024).
+We go further, however, and show how a conventional MESI-like coherence protocol can be exploited in novel ways by intelligent devices interacting with an unmodified CPU using coherence messages.
 By relaxing traditional coherence protocol assumptions (for example, the independence of cache lines), we can achieve dramatically more efficient communication between a CPU and a device.
-In the following section, we provide background and motivation for this work, including the factors that led to the descriptor-based DMA model of I/O, and why it is time to question the underlying assumptions in that model.
-In Section 3 we describe and motivate the experimental hardware platform we use for our evaluation, and also provide calibrating performance comparisons with PC data center servers.
-We show how to use PIO with PCIe in Section 4.
-In Section 5, we discuss why true cache-coherent device interconnects are fundamentally different from today’s PCIe platforms, and present a set of efficient protocols for passing messages with low latency between software running on an unmodified CPU, and a smart device which has low-level access to the coherence protocol.
-In Section 5, we go on to present the implementation of some of these protocols on our experimental platform.
-Section 6 then compares a traditional DMA-based approach with PIO over PCIe and our new coherence-based protocols using three different use-cases: (1) lightweight, synchronous local invocation of functionality on a computational accelerator, (2) I/O to and from an intelligent network interface, and (3) hardware offload of a distributed dataflow stream processor to an FPGA.
+A coherence-based message protocol which avoids PCIe significantly outperforms both DMA and PIO over PCIe and completely eliminates tail latency, while delivering the same or better throughput.
+We first provide background and motivation, including the origins of descriptor-based DMA, and why it is time to question these underlying assumptions.
+In Section 3 we describe and motivate the experimental hardware platform we use for evaluation, calibrating performance against conventional PC servers.
+We also decribe PIO with PCIe here.
+In Section 4, we discuss why true cache-coherent device interconnects like CXL3.0 are fundamentally different from today’s PCIe, and present a family of protocols that pass messages with low latency between software on an unmodified CPU, and a smart device with message-level access to the coherence protocol.
+We then present the implementation of these protocols on the hardware platform.
+Section 5 compares traditional DMA-based I/O with PIO over PCIe and our new coherence-based protocols using three use-cases: (1) lightweight, synchronous local invocation of functionality on a computational accelerator, (2) I/O to and from an intelligent network interface, and (3) hardware offload of stream processor operators.
 We survey related work in Section 6 and conclude with Section 7.
@@
 Interconnects and devices:
-In early machines, the CPU interacted with peripheral devices exclusively using programmed I/O (PIO): the device exposed hardware registers mapped into the processor’s address space.
-Both data transfer and control of the device involved reads and writes by the CPU to these registers.
-This model is also known as memory-mapped I/O (MMIO).
-This model made sense when devices were extremely simple, particularly when augmented by the use of device interrupts to obviate the need for the CPU to poll status registers on the device waiting for an I/O operation to complete.
-However, the model still requires a synchronous rendezvous between the CPU and device, and moreover requires all data to be transferred via CPU registers (since the CPU is copying data between device registers and main memory).
-DMA removes this limitation by allowing the device itself or a dedicated DMA controller to copy the data while the CPU continues to execute other code, providing parallelism and partly decoupling the CPU and device.
-Further decoupling is provided by descriptor rings: producer/consumer buffers of I/O requests in memory represented as a queue of operations, and read and written by both the CPU and the DMA-capable device.
+In early machines, the CPU interacted with peripherals exclusively using Programmed I/O (PIO): the device exposed MMIO registers, and the processor issued loads and stores to directly to these registers to perform both data transfer and control of the device.
+This model made sense when devices were extremely simple, particularly combined with device interrupts to obviate the need for the CPU to poll status registers on the device waiting for an I/O operation to complete.
+However, it still requires synchronous rendezvous between CPU and device, and moreover requires all data to be transferred via CPU registers (since the CPU is copying data between device registers and main memory).
+DMA removed this limitation by allowing the device itself or a dedicated DMA controller to copy data while the CPU continues to execute other code, providing parallelism and partly decoupling CPU and device.
+Further decoupling is provided by descriptor rings: producer/consumer buffers of I/O requests in memory, read and written by both CPU and DMA-capable device.
 In this case, interrupts need only be used when the queue of descriptors becomes full or empty.
-Almost every high-speed device today uses essentially this technique, with minor variations such as how the descriptor queues are formatted, where they are stored, and whether the head and tail of the queue is identified by registers or flags in the descriptors themselves.
+Almost every high-speed device today uses this technique, with minor variations such as the format of descriptor queues, where they are stored, and whether the queue head and tail are identified by registers or flags in the descriptors.
 This model is optimized for throughput – it is most efficient when a large volume of data must be transferred.
@@
 […]
-For the large batch transfers used in GPU-derived programming models (e.g., machine learning workloads), this latency penalty is insignificant when amortized over a large transfer.
-Indeed, such software has arguably adapted to this model imposed by the hardware (Hooker, 2021).
-Recent trends in accelerator design, however, have led to proposed standards for cache coherent peripheral interconnects, and the related idea of extending processor cache coherence to heterogeneous devices: accelerators, networking interfaces, and storage controllers (Stuecheli et al., 2018; CCIX Consortium and others, 2024; CXL Consortium, 2024; NVIDIA, 2024; AMBA® CHI Architecture Specification, 2024).
-The more far-reaching of these proposals are fully symmetric, i.e.
-devices are first-class participants in a distributed directory-based cache coherence protocol.
+However, recent trends have led to proposed cache coherent peripheral interconnects, and the extension of processor cache coherence to heterogeneous devices: accelerators, networking interfaces, and storage (Stuecheli et al., 2018; CCIX Consortium and others, 2024; CXL Consortium, 2024; NVIDIA, 2024; AMBA® CHI Architecture Specification, 2024).
+The more far-reaching are fully symmetric: devices are full participants in a distributed directory-based cache coherence protocol, as in CXL.mem 3.0 (CXL Consortium, 2024).
 Fine-grained workloads:
@@
 While end-to-end latency is mostly network propagation time in this case, latency incurred in the end system is a good proxy for key resources (such as CPU cycles) wasted before invoking the application code, and in marshaling and sending any reply.
-All these workloads incur a significant latency penalty when using DMA for data transfer. Moreover, the assumptions which led to the common DMA descriptor-based model do not hold for these workloads.
+All these workloads incur a significant latency penalty when using DMA for data transfer.
+Moreover, the assumptions which led to the common DMA descriptor-based model do not hold for these workloads.
 For example, CPU execution and I/O were decoupled using descriptor rings because CPU cycles were precious (there being few cores), and the CPU had plenty of different tasks to work on at any time.
 This is no longer true.
-Server CPUs now have 10s or 100s of cores, and moreover when running data center applications they are often dedicated to a single application, sometimes with miscellaneous functions moved to a small subset of the cores (Marty et al., 2019).
-Most of the CPU cores have nothing to do except handle small messages coming from the network.
+Modern servers now have 100s of cores (Ampere Computing, 2024) and when running data center applications they are often dedicated to a single application, sometimes with miscellaneous functions moved to a small subset (Marty et al., 2019).
+Most of the cores have nothing to do except handle small messages coming from the network.
 As a second example: for throughput-oriented workloads DMA has evolved to efficiently transfer data to and from main memory without polluting the CPU cache.
 However, for small, fine-grained interactions, it is important that almost all the data gets into the right CPU cache as quickly as possible.
-These reasons led us to rethink the DMA-based model, and whether directly involving the CPU in data transfers to and from devices might be a better approach for these scenarios on modern hardware.
-We are not the first to suggest this direction for some cases (Larsen and Ben Lee, 2015; Chaudhari et al., 2017), but we push the question much further: given a future cache-coherent interconnect, what new ways of using Programmed I/O might become compelling for data center applications?
-Note that we are not proposing abandoning the traditional model, or some of its enhancements we survey in Section 7, for throughput-oriented applications, which can continue to ride the current hardware curve for bandwidth.
+All this led us to rethink the DMA-based model, and whether directly involving the CPU in data transfers to and from devices might be a better approach for these scenarios on modern hardware.
+We are not the first to suggest this direction for some cases (Larsen and Ben Lee, 2015; Chaudhari et al., 2017; Liew et al., 2025; Schuh et al., 2024), but we push the question much further: given a future cache-coherent interconnect, what new ways of using PIO might be compelling for data center applications?
+We are not proposing abandoning the DMA model, or some of its enhancements we survey in Section 6, for large-transfer throughput-oriented applications, which can keep riding the current hardware bandwidth curve.
 Instead, we argue it is time to examine complementary alternatives.
+Figure 1. PCIe XDMA invocation latency comparison.
+Figure 2. PCIe PIO invocation latency comparison.
 3. Experimental platform
-Figure 1. XDMA invocation latency comparison.
-Figure 2. PIO invocation latency comparison.
 We investigate the modern trade-offs between PIO, DMA, and cache-coherent interconnects using a real hardware platform rather than a simulator.
@@
 Both nodes have PCIe interfaces: the ThunderX-1 has a PCIe Gen.3 x8 interface, while the FPGA has a PCIe Gen.3 x16 interface.
-In this work we connect these two interfaces with a loopback cable; this essentially replicates the case of an PCIe accelerator card and allows us to compare PCIe with ECI in our experiments.
+In this work we connect these two interfaces together with a cable; this replicates a PCIe accelerator card and allows us to compare PCIe with ECI in our experiments.
 While Enzian provides a real hardware platform with server-class performance, direct comparison with modern server platforms is not straightforward: the latter do not offer full cache coherence between CPU and devices, but have more recent, faster cores, memory systems, and PCIe.
-We therefore calibrate against a modern PC for both PIO and DMA over PCIe with benchmark comparisons of Enzian and a modern PC (Intel Core i7-7700 3.6GHz Kaby Lake, PCIe Gen.3 x16) with 64-byte cache lines connected to an AMD Virtex UltraScale+ VCU118 card, using the “write then read” experiment detailed in Section 6.1.
+We therefore calibrate both PIO and DMA over PCIe with benchmark comparisons of Enzian and a modern PC (Intel Core i7-7700 3.6GHz Kaby Lake, PCIe Gen.3 x16) with 64-byte cache lines connected to an AMD Virtex UltraScale+ VCU118 card, using the “write then read” experiment detailed in Section 5.1.
 The VCU118 uses the same FPGA as Enzian, albeit a slightly slower speed gauge.
@@
 Figure 1 shows results for various data transfer sizes, running the Xilinx driver in both interrupt-driven and polled modes.
-We see that XDMA over PCIe is about 3 times faster on the PC than on Enzian, while the difference between interrupt-driven and polling performance is much insignificant.
+XDMA over PCIe is about 3 times faster on the PC than on Enzian, while the difference between interrupt-driven and polling performance is much less significant.
 Single transaction latency is almost constant on both platforms up to the PCIe transaction size limit of 4KiB, and then increases.
@@
 In addition, the memory system is somewhat faster on the PC.
-The factor of 2 difference in PCIe bandwidth between the two platforms does not appear to be a factor in these experiments.
+The factor of 2 difference in bandwidth between the two platforms (PCIe Gen3 x8 vs.
+x16) does not appear to be a factor in these experiments.
 PIO performance over PCIe
 In the second calibration experiment the CPU performs PIO reads and writes over PCIe to the FPGA.
-We experimented with several different optimizations in this experiment (see Section 4), in particular the use of write combining and vector instructions.
-We present the best figures obtainable on each platform.
-Figure 2 shows that for transaction sizes above 32 bytes, the PC is about twice as fast as Enzian over the PCIe cable between CPU and FPGA.
-Here, the difference in PCIe bandwidth is responsible for the latency difference, in particular for reads (limited to 128 bits on both platforms).
-Writes, in contrast, are combined and pipelined at the PCIe interface.
-4. Programmed I/O over PCIe
 PCIe is complex, and we spent considerable time optimizing CPU-initiated reads and writes for different platforms.
 We describe here how we arrived at optimally-efficient PIO latency to the FPGA over PCIe.
-We pre-map PCIe apertures (BARs) into user space on the CPU and measure latency as the time to write a value to this space (thereby invoking a function on the device) and read a result back (Figure 3).
+We pre-map PCIe apertures (BARs) into user space on the CPU and measure latency as the time to write a value to this space (thereby invoking a function on the device) and read a result back (Figure 3) – this is basically uncached MMIO.
 // map 4KiB of PCIe memory for read/write
@@
 __sync_synchronize();
-Figure 3. An invocation using PIO & PCIe. Error handling omitted.
+Figure 3. Invoking with PIO & PCIe; error handling omitted.
 PCIe memory writes are posted transactions and may complete out-of-order, allowing multiple requests in-flight at the same time.
@@
 The Intel Data Streaming Accelerator (Intel, 2024) offers MOVDIR64B and ENQCMD instructions, allowing 64-byte PIO writes with write-combining mappings and user-space descriptor submission without driver intervention.
-Mapping memory with write-through attributes, as suggested by Tide (Humphries et al., 2024), allows cache line-sized PIO reads.
+Mapping memory with write-through attributes (as in Tide (Humphries et al., 2024)), allows cache line-sized PIO reads.
 These architecture-specific extensions are not available on ThunderX-1.
+Figure 2 shows the best figures we could obtain on both platforms.
+For transaction sizes above 32 bytes, the PC is about twice as fast as Enzian over the PCIe cable between CPU and FPGA.
+Here, the difference in PCIe bandwidth is responsible for the latency difference, in particular for reads (limited to 128 bits on both platforms).
+Writes, in contrast, are combined and pipelined at the PCIe interface.
+Note this benchmark permits reordering of writes, and therefore already reflects the performance achievable using the reorder unit proposed in  (Liew et al., 2025).
 4. PIO over a coherent interconnect
-We now turn to programmed I/O over a cache coherent interconnect.
-A naive approach would have the CPU simply copy data over a coherent interconnect as in conventional PIO.
-This requires careful cache management on the CPU using explicit cache flush and invalidate instructions, but is appropriate when the hardware standard mandates a cache on the device, as with OpenCAPI 4.0 (Stuecheli et al., 2018; Allison et al., 2020).
-However, we are more interested in how a coherent interconnect can best perform low-latency, efficient data movement between CPU registers and device hardware, including cases where the conventional notion of the cache as a view over memory contents is not appropriate, such as data arriving over a network.
-Scoping:
-We scope the scenario further to the case where the CPU cores, or their Last-Level Cache (LLC), communicate with devices via a MESI-like (Nagarajan et al., 2020), message-based coherence protocol with stable states corresponding to Modified (dirty), Shared (clean), Exclusive (clean), and Invalid, using a distributed cache directory.
-Note that this means that each address or cache line has a home node, which maintains the directory entry for the line.
-We also assume a symmetric protocol: each device is a first-class participant in the protocol with a subset of physical addresses homed at it, and the device therefore maintains directory state for these lines (although it might not actually cache data itself).
-Moreover, we take the perspective of the device designer: the device sees individual cache messages, rather than transparently coherent memory.
-These assumptions are rare in hardware today but they do hold for proposed standards like CXL 3.0 (CXL Consortium, 2024), and also for the Enzian platform we use in experiments.
-Baseline: simple coherence
-We first discuss using simple cache coherence to re-apply existing techniques inter-core communication in software, such as FastForward (Giacomoni et al., 2008).
-Here (Figure 4), the sender (CPU) writes data to a cache line on which the receiver is spinning reading.
-Assume the line starts in Shared state in both caches.
-When the sender writes the line (typically atomically via the core’s write buffer), it transitions briefly to Modified state, before being fetched by the receiver over the interconnect and reverting to Shared.
-Figure 4. Simple inter-core communication over coherence
-This can be highly efficient in software, particularly when deployed as a ring buffer of cache lines, and results in minimal interconnect traffic (since the receiver spins on a local copy of the line), although the CPU itself must spin.
-Indeed, this is the technique recently used in CC-NIC (Schuh et al., 2024) to emulate a NIC using a different CPU socket.
-However, this design assumes that each endpoint can only manipulate its local cache using software instructions (loads, stores, etc.) and also that flush or invalidate instructions should be avoided for performance reasons.
-This is not the case when one of the endpoints is a device with message-level access to the cache coherence protocol, and possibly maintaining its own directory state.
+While PIO over PCIe outperforms DMA on our experimental hardware over a range of payload sizes (consistent with the findings of other authors), the main contribution of this paper to show the further benefits of using a cache-coherent interconnect.
+We present a family of message protocols, all of which run over an existing MESI-like coherence protocol, which substantially outperform conventional PIO and DMA.
+Figure 4. FastForward-style coherent messaging
+Fast CPU–CPU message passing
+Our starting point is a set of software protocols used to communicate over shared memory between CPUs in a cache-coherent system, including FastForward (Giacomoni et al., 2008) and the NIC emulation used in CC-NIC (Schuh et al., 2024), and exemplified in Figure 4.
+These target machines with directory-based coherence which supports direct cache-to-cache forwarding of lines without writebacks, e.g.
+MOESI (Nagarajan et al., 2020).
+In these protocols, a receiving core generally spins reading from a cache line (usually part of a larger array organized as a ring buffer), holding it in cache in Shared state.
+The sending core starts to write to the line, causing it to be fetched in Exclusive state and invalidated in the receiver’s cache.
+When the receiving core next polls the cache line, it is fetched from the sending core and arrives back in Shared state on the receiver, where it is immediately loaded into CPU registers.
+In practice, the sender generally writes the entire line at once (exploiting a write buffer), so that in the fast case the line is transferred in two round-trips over the interconnect, as shown in Figure 4.
+Techniques like including a “finished” flag in the cache line data can be used to handle cases where the line is transferred to the receiver before it has been completely written.
+These protocols have excellent performance, and on modern hardware achieve register-to-register latencies in the low hundreds of nanoseconds.
+They are easily parallelized to many cache lines (with a single polled flag for many lines of payload, plus the necessary barriers) to achieve excellent throughput.
+This basic technique is used, for example, CC-NIC (Schuh et al., 2024) to emulate a NIC using a different CPU socket.
+However, they have several limitations.
+The first is that the CPUs must busy-wait polling memory.
+This does not generate interconnect traffic since the line being polled is in local cache, but does consume energy.
+Moreover, optimal performance depends on timing, and if the receiver polls too frequently, the line can be forwarded multiple times before it is complete, leading to wasted interconnect round-trips and increased latency.
+For these reasons, these protocols are rarely used for real-world Inter-Process Communication (IPC), even when it makes sense to dedicate cores to individual clients and servers for long periods.
 The implications of message-level access
-A device on a symmetric coherent interconnect can see more cache state and events, and also initiate more operations, than a CPU core can.
-Note that this does not entail a cache per se on the device at all, indeed this is rarely appropriate.
-First, the device receives messages from the CPU cache that request lines in Shared or Exclusive state, or which request that the caching state on the device be downgraded (e.g., from Exclusive to Shared, or from Shared to Invalid).
+The key insight behind this paper is that a device on a symmetric coherent interconnect sees more cache state and events than are visible to software, and can also initiate more operations than a CPU core can.
+Moreover, this does not require a cache on the device at all, indeed this is rarely appropriate: it is sufficient to generate and respond to individual coherence messages.
+We explain this in more detail.
+First, the device now receives messages from the CPU cache that request lines in Shared or Exclusive state, or which request that the caching state on the device be downgraded (e.g., from Exclusive to Shared, or from Shared to Invalid).
 It is free to react internally to these events as it chooses, as long as it does not violate the protocol.
-Second, the device can issue such requests itself at any time, allowing more fine-grained control of the state of a line in the CPU’s cache.
+Second, the device can also issue such requests itself at any time, allowing more fine-grained control of the state of a line in the CPU’s cache.
 Third, for lines which are homed at the device, the device’ directory maintains explicit information about the cache line state at the CPU and other nodes as well as locally.
@@
 This last insight is important because it, in combination with the previous observation, it allows us to construct efficient communication mechanisms which rely on a combination of the device’s richer view of the protocol and the software on the CPU adhering to additional conventions to communicate channel semantics to the device.
-Put simply, requests to specific cache lines can be used to signal particular higher-level operations to the device, much as register reads and writes to traditional devices frequently have semantics very different from memory access.
-This is an oversimplification, and we discuss the subtleties that arise below.
-First, however, we present four related communication protocols, layered above MESI-like coherence, which provide for efficient, blocking communication between a conventional CPU and a coherence-aware device.
-(a) CPU-initiated, device-homed
-(b) Device-initiated and homed
-(c) CPU-initiated, CPU-homed
-(d) Dev.-initiated, CPU-homed
+This means requests to specific cache lines can be used to signal particular higher-level operations to the device, much as register access to traditional devices frequently has very different semantics to memory access.
+(a) Write to device
+(b) Read from device
+(c) Read and write combined
 Figure 5. Protocol variants for efficient CPU-device messaging
-Communication protocol design variants
-We present 4 different but similar communication protocols (shown in Figure 5) which vary based on whether (a) the CPU or the device initiates the message transfer, and (b) the cache lines used are homed on the CPU or device.
-All provide a synchronous invocation: a transfer of a single cache-line-sized message, followed by a single cache-line response.
-All use 2 cache lines, which we designate A and B.
-These protocols can be generalized to larger payloads, and latency will grow linearly with the payload size.
-Interestingly, they can be made slightly more efficient by using more cache lines, but quickly become complex.
-We focus on the simple 2-line case in this paper.
-The protocols use the same techniques and insights, and are best described using the example of a CPU-initiated invocation on the accelerator using addresses homed on the device (5(a)).
-An implementation of this protocol is used for the evaluation in Section 6.1.
-Note that, unlike RPC, invocations are local to the machine and avoid marshaling or unmarshaling of messages.
+CPU-device message passing with coherence
+We can now present a family of related point-to-point messaging protocols which can be built over a symmetric, directory-based cache coherence protocol where one end-point (the device) has direct access to the low-level cache messages.
+Figure 5 shows three such protocols, all with the same broad structure, operating on a pair of cache lines which are homed at the device.
+This allows the device to precisely track states of the cache lines used and greatly simplifies the protocol.
+They vary according to whether data flows (a) from CPU to device, (b) from device to CPU, or (c) both ways.
+Time is on the vertical axis.
+Under each end-point (CPU or device) are two columns, one for each cache line, giving its state over time.
+The notation E->M within a cell indicates the silent upgrade of a line from Exclusive to Modified, while the arrows between cells indicate a state changes corresponding to visible interconnect messages.
+The first and second protocols implement uni-directional message passing between the CPU and device, and are the basis for the Ethernet NIC presented in Section 5.2.
+The third option combines the two by packing a request and response into one invocation of the protocol (i.e.
+2 round-trips) and is best suited for send-and-receive interactions like RPC.
+We demonstrate this in Section 5.1.
+Each protocol requires just two interconnect round trips, beginning with the two lines in a defined quiescent state.
+At each point one of the lines (which we refer to as $A$) serves to transfer information from CPU to device, and the other ($B$) from device to CPU.
+Over the course of one transaction both lines migrate from one cache to the other, ending in the same (caching protocol) state in which the other began.
+The two then swap roles for the next iteration.
+There is no actual cache on the device: any payload is delivered directly into the device’s execution pipeline, or read from an output register, as appropriate.
+The following describes the execution of variant (c) in detail:
   1. (1)
-    The protocol starts with line B in Exclusive state in the CPU cache, and line A Invalid (we show how this can be conveniently achieved below).
-To initiate an operation, the CPU writes the request data from its registers into line B, which transitions to Modified.
-This happens without communication with the device, since B was previously in Exclusive.
+    $B$ starts Exclusive in the CPU cache, and $A$ Invalid, with the (directory) states the opposite on the device.
+Software writes payload data from registers into $B$, which (silently) transitions to Modified.
   2. (2)
-    To signal to the device that line B holds a new request, the CPU then executes a load to line A.
-Since A is Invalid, the cache requests A in Shared from the device.
-This violates the typical convention that cache line operations are independent, but the device is free to interpret this request for A as a signal from the CPU that B holds a request.
+    To initiate the exchange, software reads from $A$, triggering a load shared message to the device. The device interprets this as a signal that $B$ contains fresh data.
+    This coupling of independent line states is the first deviation from ‘standard’ coherence, and is impossible to achieve purely in software. It is nevertheless entirely invisible to the CPU.
   3. (3)
-    The device does not respond immediately to the request for A, thereby stalling the load and blocking the requesting core.
-Instead, it requests B from the CPU’s cache, in Exclusive.
-The interconnect therefore transfers the request data to the device, and line B duly transitions to Invalid in the CPU’s cache.
+    The device now stalls the CPU by not responding to the load of $A$ until it has data to send. Instead, it requests $B$ in Exclusive to fetch the payload, leaving it Invalid in the CPU’s cache.
   4. (4)
-    The device reads the data it has received in B, and computes a result on it, which is to be returned to the CPU. It returns this using the response to the CPU’s request for cache line A.
+    The device uses the data in $B$ to computes a result which it returns as a result payload when it finally responds to the CPU’s load from $A$.
   5. (5)
-    The CPU cache receives line A’s new data and the CPU unblocks, immediately loading the first word of the invocation response into a register.
-As described, the end state is that A is now Invalid and B is Shared in the CPU cache.
-However, in some deployed MESI-like protocols (including the Enzian interconnect we use), the device is free to return the data stipulating that it is not Shared but now Exclusive to the CPU.
-This final feature leaves the cache in the state it started in, with the roles of A and B reversed, ready for the next transaction.
-The whole exchange takes 2 round trips on the coherent interconnect, and require no spinning on the CPU.
-The other 3 protocols work similarly.
-For example, 5(b) shows the analogous case where the device initiates the request to the CPU, but still using lines homed on the device.
-This is used for the NIC experiment in Section 6.3.
-The remaining two cases, where the cache lines are homed on the CPU, are more complex to deal with, since there is a possibility that the CPU cache will evict one of the cache lines without warning.
-The device will not observe this, and so might incur an access to main memory on the CPU node.
-In this paper we therefore focus on the case where the cache lines are homed on the device.
-Subtleties:
-A number of issues must be handled to make this practical, starting with timeouts: we block a core’s request for a cache line until a device operation has completed.
-A timeout in the CPU’s cache is likely to cause a memory fault exception, a fatal machine check, or the processor simply locking up (on Enzian, it is a machine check).
+    The CPU cache receives $A$ and the CPU unblocks, immediately loading the first word of the response into a register.
+This would lead to $A$ in Shared in both nodes, since software requested $A$ using a load.
+However, as an optimization, the device can return the line in Exclusive instead, invalidating its own copy, so the protocol returns to a quiescent state with the roles of the lines reversed.
+This protocol either eliminates or minimizes the drawbacks of a software-only protocol: the ability to stall the CPU eliminates both the need to poll, and the race condition.
+By coupling and overlapping two coherence transactions we are able to transfer data in both directions with the same two round-trips.
+All data is transferred atomically so there is no risk of half a payload being shipped.
+This is the best achievable result with an unmodified coherence protocol, since no cache transaction both sends and receives a payload.
+Figure 6. Invocation latency over ECI vs. FastForward
+Figure 6 shows the distribution of total latency for this operation implemented on Enzian.
+Median latency is around 1600ns when the result is returned in Shared state as requested by the CPU (“ECI unopt”).
+When we return the line in Exclusive, this drops to around 900ns (“ECI”).
+For comparison, we also show an implementation of the FastForward protocol (Giacomoni et al., 2008) exchanging cache lines between CPU sockets in a dual-socket ThunderX-1-based Gigabyte R150-T61 server with similar CPU and DDR specification to Enzian; this achieves a median latency of about 1750ns.
+The optimized protocol is performing 2 round-trip message exchanges over ECI, which has a one-way latency at the link layer of about 150ns.
+The rest of the overhead (300ns) is mostly incurred in the protocol processing in the directory controller, and would be lower in an ASIC implementation.
+The FPGA is clocked at about 300MHz.
+Returning a line in Exclusive
+The ability for the device to unilaterally return a line in Exclusive delivers an important performance benefit.
+The ECI protocol, and the ThunderX-1 last-level cache, both support this.
+Interestingly, it is also proposed in the (as yet unimplemented) CXL.mem 3.0 standard (CXL Consortium, 2024), along with the “back invalidate” operation providing symmetric cache coherence.
+We suggest this is a useful feature for future protocols.
+Handling larger messages
+The protocols above handle one line at a time (128B on Enzian).
+They can be used repeatedly to transfer more data than this, but this is inefficient.
+Instead, we extend the protocols described so far using parallelism and pipelining to give much better throughput.
+This leads to two further protocols, which are the ones we evaluate in Section 5.
+For the NIC example of Section 5.2, with predominantly unidirectional communication, we dedicate one pair of cache lines as control (as in 5(a) and 5(b)) and add overflow cache lines to carry data.
+When the protocol invalidates the control lines, it also invalidates these overflow cache lines in parallel, exploiting the full bandwidth of the coherent link.
+We use this technique to transfer packets larger than one cache line, up to an Ethernet jumbo frame (9kB).
+We can do even better when the application has a symmetrical communication pattern, i.e.
+the CPU sends and receives similar amounts of data in one invocation.
+This is the case for the device invocation and Timely Dataflow experiments we show in Section 5.1 and Section 5.3.
+In this case, the better option is to prefetch multiple instances of the protocol shown in 5(c).
+To transfer $n$ cache lines at a time, we maintain two groups of $n$ lines.
+The CPU first writes all arguments for the device to the lines in group 0, and then prefetches all cache lines in group 1.
+The prefetches trigger the device logic to invalidate all cache lines in group 0, fetching the arguments.
+The prefetches are issued in parallel by the CPU and again provide enough concurrency to saturate the link.
+Handling timeouts
+The protocols block a core’s request for a cache line until a device operation has completed.
+This timeout is typically quite long (often hundreds of milliseconds) but if it happens the CPU’s cache is likely to cause an exception, or other fault like a machine check.
 We solve this issue with a small state machine on the device that returns a “not ready yet” response before an impending timeout, causing the software on the issue another request for the other cache line, extending the response time indefinitely without the need for spinning.
 In practice, we target operations that are shorter than the timeout.
-The second issue concerns deadlock due to the device stalling a request until it can issue, and get a reply back for, another request to a different cache line.
-Existing standards for interoperable cache coherent interconnects are silent on what happens in this situation, possibly because it has not occurred until now.
+Avoiding deadlocks
+It is also important to prevent deadlocks due to the device stalling a request until it can issue, and get a reply back for, another request to a different cache line.
+Existing standards for inter-operable cache coherent interconnects are silent on what happens in this situation, possibly because it has not occurred until now.
 The design assumes that transactions on different cache lines can progress independently.
-In most implementations, this assumption typically holds in order to maximize the memory bandwidth utilization and minimize request latency, and also to simplify reasoning about deadlock freedom of the coherence protocol itself.
-However, it may be the case that the cache stripes transactions across a limited number of independent units and which might deadlock if both A and B were mapped to the same unit.
+In some implementations, this assumption holds in order to maximize the memory bandwidth utilization and minimize request latency, and also to simplify reasoning about deadlock freedom of the coherence protocol itself.
+However, it may be the case that the cache stripes transactions across a limited number of independent units which might deadlock if both $A$ and $B$ were mapped to the same unit.
 Were this to be the case, it could be avoided by careful placement of $A$ and $B$ in the physical address space.
-This is actually the case in our FPGA device implementation, described in Section 5.
-Implementation in Enzian
-We implemented the two device-homed protocols over Enzian and evaluate them in Section 6.
-The implementations use SystemVerilog and SpinalHDL (contributors, 2024) and exploit an existing “Directory Controller” on Enzian’s FPGA (Ramdas, 2023) which manages the coherent links to the CPU and handles transient states in the ECI protocol.
-ECI provides all the messages we require to implement our messaging protocols.
-In particular, it includes the ability for the device to return a line to the CPU cache in Exclusive state, even though the CPU requested it in Shared.
-This optimization dramatically reduces latency, as shown in Section 6.1.
-In our case, no flow-control issues arise since ECI uses separate virtual channels for request and reply messages, removing the chance of deadlock.
-However, other complications arise: ThunderX-1 generates Global Synchronization (GSync) messages which function as global barriers.
-A GSync blocks any subsequent transactions until all those started before the GSync have retired.
-Fortunately, any invalidations sent are not blocked by core, allowing the protocol to progress regardless.
+This is actually the case in the Enzian FPGA directory controller implementation we use (Ramdas, 2023), which follows similar provisioning choices in the CPU.
 The ThunderX-1 divides its last-level cache functionality into a set of units called TADs, each of which can handle up to 16 simultaneous cache transactions.
 We ensure that consecutive cache lines are mapped to different TADs to ensure that the operations can proceed independently.
+Another possible source of deadlocks is that the CPU and L2 cache might issue requests out of order, especially in the case of using prefetches.
+This is mainly relevant for handling requests larger than one cache line, as we discussed above.
+The device implementation needs to be aware of possible reorderings and should not rely on the request ordering to perform state transition to avoid deadlocks.
+In our implementation, we always buffer possible in-flight cache lines on the FPGA and advance state machines based on number of requests we see, without relying on the actual order of them.
+Correctness and interoperability
+An obvious concern given the complexity of underlying cache coherence protocols is the correctness of the implementation.
+An error in the device might cause the state machine in the CPU’s cache to enter a state from which it cannot recover.
+Worse, while our implementation relies on an existing, and well-tested, directory controller implementation on the FPGA, we are using the ECI protocol in unconventional ways unforeseen by the designers of the CPU and its cache.
+In practice, after extensive testing we have not seen protocol-related problems even under heavy load.
+Moreover, we are reassured by recent work on exhaustive semi-open-state testing of coherence protocols (Schult et al., 2024) which can formally define the state space of a subset of a coherence protocol (which can nevertheless include the kinds of behaviors we describe here), and then exhaustively test a hardware implementation against that formal model.
+Enzian-specific implementation issues
+Our implementations will all be made available as open-source and use a combination of SystemVerilog and SpinalHDL (contributors, 2024), and use a version of the existing Enzian FPGA directory controller (Ramdas, 2023).
 The ARMv8-A weak memory model means that barriers are needed to reliably order reads and writes.
@@
 On the ThunderX-1 this is sufficient to ensure ordering observed at the FPGA.
+Generality
+If we had only simulated the protocols in Figure 5 rather than implement them in a hardware platform, it seems unlikely that we could have become aware of the numerous subtleties and challenges that real coherent interconnects present.
+However, this raises generality concerns, since our implementation is tied to a single platform.
+In principle, all the protocols we have described here could be implemented over any coherent interconnect which uses a symmetric, MOESI-like, distributed directory-based model of coherence, and where the interconnect hardware exposed major state transitions to the device-specific logic which served as the end-point for messaging.
+Several current and future standards satisfy these conditions, for example CCIX (CCIX Consortium and others, 2024), TileLink (Berkeley Architecture Research, 2022), and CXL.mem 3.0 (CXL Consortium, 2024), although notably neither CXL 2.0 nor CXL.cache 3.0.
+We are also seeing something of a boom in custom interconnects from a range of vendors.
+By making the requirements for this fine-grained, low-latency communication clear, future designers can take this into account when designing new interconnects and parts.
 5. Evaluation
-We compare PIO over a coherent interconnect with two other reference points: PIO directly over PCIe– essentially, the CPU copying data to and from a region of PCIe address space – and descriptor-based DMA over PCIe.
-All use the same Enzian hardware platform.
-Coherent PIO uses the Enzian native ECI, while the two PCIe techniques connect the CPU and FPGA PCIe interfaces back-to-back with a loopback cable as described in Section 3, effectively turning the Enzian FPGA into a conventional PCIe FPGA card.
-DMA experiments all use the Xilinx XDMA hardware IP on the FPGA.
+We compare coherent PIO with our two other reference points: optimized PIO over PCIe described in the previous section, and descriptor-based DMA over PCIe.
+All use the same Enzian platform.
+Coherent PIO uses the Enzian native ECI implementation, while the PCIe approaches use a loopback cable between PCIe interfaces on the CPU and FPGA, effectively turning the Enzian FPGA into a conventional PCIe FPGA card.
+DMA experiments use the Xilinx XDMA IP.
 We explore three different application scenarios: synchronous invocation to an accelerator on the FPGA, a closely-coupled high-speed NIC, and a stream processor which moves some of its dataflow operator graph to the FPGA.
-6.1. Invoking functions on an accelerator
-We first measure latency when the CPU invokes a function on the device, passing arguments, and obtaining a result. We use the protocol shown in 5(a), whereby the cache lines are homed on the FPGA.
-Figure 6. Invocation latency over ECI vs. FastForward
-Figure 6 shows the distribution of total latency for the this operation.
-We see a median latency of around 1600ns for the unoptimized protocol, wherein the result line is returned in Shared state as it was requested by the CPU.
-When we introduce the optimization of returning the line in Exclusive, the median latency drops to around 900ns.
-For comparison, we also show an implementation of the FastForward protocol (Giacomoni et al., 2008) exchanging cache lines between CPU sockets in a dual-socket ThunderX-1-based Gigabyte R150-T61 server with similar CPU and DDR specification to Enzian; this achieves a median latency of about 1150ns.
-The optimized protocol is performing 2 round-trip message exchanges over ECI, which has a one-way latency at the link layer of about 150ns.
-The rest of the overhead (300ns) is mostly incurred in the protocol processing in the directory controller.
-The FPGA is clocked at about 300MHz.
-A significant penalty is incurred when the results of the call are returned in Shared, resulting in further round trips over ECI to (unnecessarily) invalidate the line at the FPGA before sending the next message.
-This suggests that future coherent interconnects would benefit greatly from allowing the home node to return a line in Exclusive.
-Overall, the optimized protocol comfortably outperforms FastForward, despite the latter exploiting highly optimized hardware implementations at both ends (clocked at 2.1GHz) and aggressively polling, whereas our optimized protocol requires no spinning on the part of the CPU or device.
-6.2. Invocation over ECI vs. PCIe PIO and DMA
+With the exception of the NIC case, the results show very little variance. All graphs show 95% confidence intervals, but these are typically not visible.
+5.1. Accelerator invocation
 Figure 7. Invocation latency for different payload sizes
-We now compare the latency achieved in the previous section with both PIO over PCIe and the Xilinx XDMA hardware using descriptors.
-For DMA numbers we use the Xilinx XDMA benchmark (Xilinx DMA Benchmark, 2024), which essentially performs the same operation using descriptor-based DMA.
-In both these cases, the FPGA is configured to map the request and response into a write to, and a read from, onboard Block RAM.
-As in Section 6.1, we measure latency to send a request from the CPU to the FPGA and read a response.
-However, this time we vary the payload size (of both request and response) and report median latency figures.
+We first explore the scenario where the CPU invokes a function on the device configured as an accelerator, using the three communication approaches to send argument data and receive result data.
+In this case, the FPGA is configured to map the invocation into a write to, followed by a read from, onboard Block RAM.
+For DMA numbers we use the Xilinx XDMA benchmark (Xilinx DMA Benchmark, 2024), which performs the operation using descriptor-based DMA.
+We implement coherent PIO with the protocol shown in 5(c) to send a request and receive a response in two round-trips.
+Latency comparison
+We first compare the latency for the CPU to complete a function invocation on the device, measured as the time taken for the CPU to send a request from the CPU to the FPGA and read a response.
+We vary the payload size (of both request and response) and report median latency figures.
 The results are shown in Figure 7.
-Up to payloads of 8KiB, the latency of XMDA-based communication is dominated by descriptor setup and manipulation, and remains largely independent of payload size.
-Faster PCIe interconnects (this is a PCIe Gen3 x8 link) would amplify this effect.
-Latency over ECI is dramatically lower for smaller payloads.
-Unsurprisingly, it is constant up to 128 bytes (the size of a cache line on Enzian) and then increases linearly, since the protocol simply replicates each cache line exchange until all the data has been transferred.
-It is only at 8KiB that ECI invocation latency exceeds that of XMDA.
-Arguably the transfer of multiple cache lines over ECI might be improved by clever pipelining and locating the lines in physical memory so as to exploit parallelism in the directory controllers, but we do not explore that in this paper.
-Relative to ECI, PIO over PCIe performs poorly except for payloads under 16 bytes in size. This is due to limitations in PCIe that we have discussed in Section 4, notably the limited read bus width.
-We conclude that up to the page size (4KiB) or the typical size of jumbo frames, PIO remains significantly faster than the DMA option commonly used for these transaction sizes.
-Thus even using the XDMA implementation for the experiments we still have the room for comparison with optimized DMA implementations.
-This experiment is closed-loop and single-threaded, and so achieved throughput is proportional to the round-trip latency.
-Fully saturating the raw link bandwidth (in all cases) would require many parallel invocations at once, which would go against our goal of supporting applications where latency is paramount.
+Up to payloads of 8KiB, the latency of DMA-based communication is dominated by descriptor setup and manipulation, and remains largely independent of payload size (upper scale).
+Faster PCIe interconnects or CXL would only amplify this effect.
+Latency over ECI is dramatically lower across the board.
+It is constant up to 256 bytes (two Enzian cache lines) and then increases slowly, but linearly.
+This is due to the pipelined nature of the ECI PIO protocol: before we saturate the links, the incremental latency induced by each extra line is small.
+Relative to ECI, PIO over PCIe performs poorly for any payload more than 16 bytes, due to the inherently-limited PIO read bandwidth discussed in Section 3, but competes well with DMA up to 2KiB payload.
+We conclude that, for almost all transfers up to and beyond 8 KiB, coherent PIO is significantly lower latency than both the DMA option commonly used for these transaction sizes, and PIO over PCIe.
+More PCIe bandwidth (e.g.
+PCIe Gen5) will not change the PIO results significantly without support for more efficient MMIO reads, since PCIe round-trip latency remains the same.
+Neither would it improve DMA latency, which is dominated by descriptor overhead.
+In contrast, the lower interconnect latency available with newer CXL versions would improve things, but would also deliver the same benefit to the coherent PIO case.
+Figure 8. Invocation throughput for different payload sizes
+Throughput comparison
+We also show throughput figures, omitting PIO over PCIe since performance (as seen above) is not practical for larger transfers.
+We measure how many back-to-back invocations one CPU core can make for 1 ms, using different transaction sizes.
+We calculate throughput by the total data volume transferred divided by the time elapsed and report the median value in Figure 8.
+ECI PIO throughput increases with transaction size up to 32KiB payloads to a peak of 2.19 GiB/s and then dropping slightly, due to thrashing in the 32KiB L1 data cache.
+It performs better than PCIe DMA over all transaction sizes presented, and by a comfortable margin.
+At 64 KiB, we are still far from saturating throughput for PCIe DMA; it is well known that very large transaction sizes are needed to achieve maximum throughput.
+In conclusion, ECI PIO shows a clear throughput advantage over PCIe DMA at multiple page sizes (up to 64KB) for device invocations, to the extent that the determining factor is less the throughput achieved and more the number of CPU cycles or cache size available.
+As with the latency, this is unlikely to change with faster PCIe interconnects, and CXL would likely result in improvements across the board, including for the coherent PIO.
+The results also show that there is an optimal native transaction size for coherent PIO, in this case the L1 cache size: larger transfers should be broken down into smaller transactions of optimal size to achieve maximum throughput.
 5.2. Network interface adapter
@@
 We implement 3 variations of a 100 Gb/s NIC in Enzian.
-The first (8(a)) is a conventional approach connecting one of Enzian’s 100 Gb/s Ethernet MACs directly to the XDMA engines, resulting in a NIC programmed using descriptor rings.
-The second design (8(b)) buffers packets in onboard SRAM on the FPGA and allows the CPU to read and write packet and control data using PCIe reads and writes.
-The final approach (8(c)) uses a custom module on the FPGA to bridge between the ECI directory controller and the Ethernet MAC, using a slight variant of the protocol in 5(b) to deliver network packets to the CPU in multiple cache lines.
-In all cases we clock the core NIC logic at 250 MHz.
-Figure 9. NIC implementations: receive latency
-Figure 10. NIC implementations: transmit latency
+The first (9(a)) is a conventional approach connecting one of Enzian’s 100 Gb/s Ethernet MACs directly to the XDMA engines, resulting in a NIC programmed using descriptor rings.
+The second design (9(b)) buffers packets in onboard SRAM on the FPGA and allows the CPU to read and write packet and control data using PCIe reads and writes.
+The final approach (9(c)) uses a module on the FPGA to bridge between the ECI directory controller and the MAC.
+A variant of the protocol in 5(b) delivers network packets to the CPU in multiple cache lines, and 5(a) is used to send a packet.
+The NIC logic is clocked at 250 MHz.
+Figure 10. Latency of NIC implementations
 Table 1. NIC implementations: latency percentiles on selected packet sizes. Higher tail latencies are marked in bold.
@@
 9600 39.43 39.48 39.50 39.55 9.07 9.19 9.65 9.95
-For experiments we configure the Ethernet MAC in near-end PCS/PMA loopback mode for a reliable measurement of packet delivery inside the server without network delays.
-We define the receive latency as time passed from the last beat of the packet appearing on the ingress streaming interface of the Ethernet MAC, to the CPU fully receiving the packet content in its registers.
-The transmit latency is defined similarly as time passed from the CPU having the packet ready in registers, to the last beat of the packet appearing on the egress streaming interface on the MAC.
-We do not include the time it takes the packet to go through the Ethernet PCS/PMA loopback, since the MAC is fixed and the same in all implementations presented in this experiment.
+For experiments we configure the Ethernet MAC in near-end PCS/PMA loopback mode to reliably measure packet delivery inside the server without network delays.
+We define receive latency as time from the last beat of the packet appearing on the ingress streaming interface of the MAC, to the CPU fully receiving the packet contents in its registers.
+Transmit latency is similarly defined as time from the CPU having the packet ready in registers, to the last beat of the packet appearing on the egress streaming interface on the MAC.
+We do not include the time for the packet to go through the Ethernet PCS/PMA loopback, since the MAC is fixed and the same in all implementations presented.
 To minimize disruptions from the Linux scheduler, we follow the common practice to free one CPU core from kernel tasks and IRQ processing (Akkan et al., 2012), with the isolcpus function together with core pinning using taskset.
@@
 We show figures for XDMA in both polled and interrupt-driven modes; in practice the difference in latency is small.
-Figure 9 shows receive latency for different packet sizes for each of the NIC implementations.
-As in the invocation experiment, DMA latency is fairly stable across all packet sizes, suggesting that the latency is dominated by various DMA overheads, such as descriptor setup and manipulation.
-Note that system calls here actually result in little overhead (only a few $\mu$s); the overhead is dominated by cache misses manipulating DMA data structures in main memory.
+Figure 10 shows receive latency for different packet sizes for each NIC implementation.
+As before, DMA latency is stable across all packet sizes, suggesting that it is dominated by various DMA overheads, such as descriptor setup and manipulation.
+Note that system calls here actually result in little overhead (only a few $\mu$s); the overhead is dominated by cache misses manipulating DMA data structures in RAM.
 In contrast, PIO over both ECI and PCIe offer much lower latency for small packet sizes up to 1024 bytes DMA, highlighting high efficiency gains in PIO compared to DMA.
-However, the latency of PCIe PIO quickly degrades for larger packets due to the non-posted read problem discussed in Section 4, reaching over 350 $\mu$s for 8 KiB packets.
-In contrast, ECI offers dramatically better latency even at 8 KiB packets.
-Figure 10 shows transmit latency.
-In this case, PCIe PIO competes well with ECI across all packet sizes due to combining posted writes as we discussed in Section 4.
-PIO over ECI shows a slightly higher latency for small packets, due to it needing 2 round-trips over ECI for each 128 bytes, versus a single PCIe round trip.
+However, the latency of PCIe PIO quickly degrades for larger packets due to the non-posted read problem discussed in Section 3, reaching over 350 $\mu$s at 8 KiB packets.
+In contrast, ECI offers dramatically better latency even at this size.
+Figure 10 also shows transmit latency.
+PCIe PIO competes well with ECI across all packet sizes due to combining posted writes.
+PIO over ECI has slightly higher latency for small packets, since it needs 2 round-trips over ECI for each 128 bytes, versus a single PCIe round trip.
 Both PIO solutions perform significantly better than DMA across all packet sizes.
-Table 1 shows the tail latency impact for selected packet sizes for each of the NIC implementations.
+Table 1 shows the tail latency impact for selected packet sizes for each NIC implementation.
 We select three representative packet sizes to evaluate tail latency: 64 for the smallest possible packet, 1536 for the normal Ethernet MTU, and 9600 for a common jumbo frame MTU.
 We use the tickless kernel setup mentioned earlier to avoid the 250 Hz Linux timer, in order to reveal the actual tail latency impact from our NIC.
-We see that PIO over PCIe results in low tail latency compared to the descriptor-based approach, while ECI completely eliminates tail latency in this scenario.
-These results show a clear advantage of using PIO over DMA, even for larger packets, and especially when using a coherent interconnect .
+PIO over PCIe results in low tail latency compared to the descriptor-based approach, while ECI completely eliminates tail latency in this scenario.
+These results show clear benefits of using PIO vs.
+DMA, even for large packets, and especially when using a coherent interconnect.
 The poor receive performance of PCIe-based PIO for receiving packets is due to the unposted-read limitation of PCIe, and disappears for transmit.
-6.4. Accelerating Timely Dataflow
-In this example we apply the techniques and results outlined so far to a practical workload: Timely Dataflow (Murray et al., 2013), a real-time stream-processing system.
-We compare the baseline performance to a partially FPGA-offloaded implementation under three CPU–FPGA communication scenarios: DMA over PCIe, PIO over PCIe, and PIO over ECI.
-Timely Dataflow schedules operators in a user-defined dataflow graph to evaluate complex operations on large streams of data, operating on variable-sized batches.
-Operators may be arbitrarily complex and are independent, and thus benefit from offloading where the overhead of shipping the data to the FPGA is compensated by greater throughput of the offloaded operators.
-To test the limits, we construct a worst-case scenario: a large (31-element) graph of simple operators (filters).
-The large graph maximizes communication overhead in the form of progress tracking data, while the simple operators leave the FPGA with only a small performance lead over the CPU.
-We want to predict whether PIO can make FPGA offload work in regimes where the fixed overheads of DMA make it impractical, for example where small batch sizes are needed to ensure data freshness (as in Differential Dataflow (McSherry et al., 2013)).
-The first (blue) bars of Figure 11 show the baseline performance of the non-offloaded (CPU-only) system against batch size.
-Note that the horizontal axis is a log scale; All four curves are actually straight lines, with offsets and slopes obtained by a least-squares fit listed in Table 2 along with the coefficient of determination ($r^{2}$).
-The remaining bars cover the three offload cases.
-Each is the median of 1000 samples.
-While no offload configuration beats the baseline, the two PIO cases outperform DMA up to batch sizes of 512B (for PCIe) and 4KiB (for ECI), with ECI almost matching the baseline at the smallest batch size (128B).
-To understand why offloading fails to produce a speedup in this example, and what changes will be necessary to allow it to, we look more closely at the fitted coefficients in Table 2.
-As shown by the $r^{2}$ values, the fit for each dataset is excellent.
-We can thus model the time taken to process a batch as a fixed component, or overhead, independent of the amount of data processed (e.g., running the Timely scheduler, or setting up DMA descriptors) plus a component which scales linearly with the size of the batch.
-Compared to the baseline, DMA offload has a significantly lower per-byte cost ($3.7$ vs.
-$8.7ns/B$) but higher fixed cost ($336\mu s$ vs.
-$204\mu s$).
-This includes the cost of descriptor setup and interrupt handling, while the per-byte cost shows the higher throughput of the FPGA.
-This includes both computation and data transfer but as DMA makes very efficient use of the (PCIe 3.0 x8) interconnect, once setup costs are paid, the per-byte transfer cost is on the order of $0.1ns$, allowing us to estimate the per-byte cost of FPGA compute at $3.6ns/B$.
-The CPU figure represents only compute.
-This linear model predicts a break-even at 26KiB.
-This matches the observed data (not plotted) where DMA offload first outperforms the baseline at a batch size of 32KiB.
-Wherever the FPGA compute cost is at least $0.1ns$ lower per byte than the CPU, DMA offload will outperform for large enough batches.
-However, due to the large fixed overhead it can never outperform the CPU for batches smaller than 16KiB.
-This limits its usefulness where small batch sizes are needed.
-While they make less efficient use of the interconnect bandwidth (as indicated by their higher per-byte costs), their lower fixed overhead allows the PIO modes to significantly outperform DMA at smaller batch sizes.
-What would it take for the faster of the two (ECI) to outperform the CPU at a moderate batch size?
-As the FPGA pipeline is identical in all cases, the limiting factor is the per-byte transfer cost.
-Taking our above estimate of FPGA compute at $3.6ns/B$, we can estimate this cost at $30-3.6=26.4ns/B$.
-Letting $S$ stand for batch size and $C$ for this transfer cost, the two break even where
-$204\mu s+S\times 8.7ns/B=221\mu s+S\times(C+3.6ns/B)$
-Rearranging gives
-$C=5.1ns/B-17\mu s/S$
-5.1ns/B being the speedup due to FPGA processing and $17\mu s$ the fixed overhead relative the baseline, the negative sign due to the higher-than-baseline overhead.
-From this we see that given the overhead the smallest batch for which a speedup is possible is around 3300B, and that a batch size of 4096B demands a transfer cost not more than 0.95ns/B, or 120ns per 128B cache line.
-As the existing protocol serializes on each line, this cost includes the round-trip latency.
-As the hardware can handle many in-flight transfers at once, modifying the protocol to transfer 32 lines at once would reduce the per-byte cost enough that ECI-PIO offload breaks even.
-Additionally, bringing the fixed cost down reduces the smallest practical batch.
-If the relative overhead could be brought below 500ns/batch then with parallel cache lines giving a 1ns/B transfer cost, FPGA offload using ECI-PIO would outperform the CPU from a batch size of 128B.
-With such an optimized protocol, ECI-PIO would outperform all other cases until a batch size of 142KiB, above which PCIe-DMA would overtake it.
-Figure 11. Timely Dataflow offload, graph size 31
-Table 2. 1st-order (offset + slope) coefficients fitted to the data series of Figure 11, as cost per byte and fixed overhead
-Time/Byte (ns) Overhead ($\mu$s) $r^{2}$
-Baseline 8.7 204 1.00
-PCIe-DMA 3.7 336 0.99
-PCIe-PIO 88 277 1.00
-ECI-PIO 30 221 1.00
-6.5. Offloading Bloom Filters
-In a second acceleration experiment we offloaded Bloom filters (Bloom, 1970) to the FPGA.
-Bloom filters are used to efficiently check whether an element is a member of a set.
-It uses little memory and allows for fast lookups.
-A Bloom filter is equipped with $k$ different hash functions and a lookup table.
+This advantage should translate to future CXL3.0 and PCIe Gen6 NICs.
+5.3. Offloading Timely Dataflow
+Finally, we look at hardware offload of operators in a real-time stream processing system: Timely Dataflow (Murray et al., 2013; McSherry, 2025).
+Timely schedules operators in a user-defined dataflow graph to evaluate complex functions over on data streams, operating in variable-sized batches.
+Operators are independent and may be arbitrarily complex, thus benefiting from hardware offloading where the performance improvement outweighs the overhead of shipping data to the accelerator.
+Timely also illustrates the flexibility of the ECI PIO protocol: our offloading implementation the protocol in Section 4 synchronously to exchange statistics between FPGA and CPU before and after processing a batch, and operators asynchronously for processing data in the batch in parallel.
+We modify the latest Rust implementation of Timely (McSherry, 2025) to include the ability to partition its dataflow graph between software operators and those implemented on the FPGA, with automatic insertion of communication channels between the CPU and FPGA to transfer both data and progress tracking information across the hardware/software boundary.
+We perform two experiments: an extreme synthetic benchmark with little actual computation, and real operators in the form of offloaded Bloom filters.
+Communication-only graph
+The synthetic benchmark uses 31-element pipeline graph of trivial filter operators (filters).
+This large operator graph maximizes communication overhead in the form of Timely progress tracking data, while the simple operators mean the FPGA has little performance advantage (if any) over the CPU, mimicking cases where small batch sizes are need to ensure data freshness, as in Differential Dataflow (McSherry et al., 2013).
+Figure 11. Timely synthetic filter graph offload
+The first (blue) bars of Figure 11 show the baseline performance of the non-offloaded (CPU-only) system against batch size. The remaining bars cover the three offload cases.
+ECI PIO batch latency is lower than both PIO and DMA over PCIe for all batch sizes, even a single cache line (128B).
+Moreover, PIO over ECI is the only technique that delivers lower latency than the software-only Rust implementation, even in this near-worst-case scenario for hardware offload.
+Bloom filters
+In our second Timely experiment we offload Bloom filters (Bloom, 1970) to the FPGA, replicating the application scenario used to evaluate Fleet (Thomas et al., 2020).
+Bloom filters are used to efficiently check whether an element is a member of a set, using little memory.
+A Bloom filter consists of $k$ different hash functions and a lookup table.
 Testing the presence of an element requires computing these $k$ hash functions and querying the lookup table with the results.
-We took inspiration from the applications discussed in Fleet (Thomas et al., 2020).
-As the CPU baseline, we re-implemented the Fleet C++ code in Rust as a Timely Dataflow operator.
-In the Fleet implementation, the hashes were computed in parallel for each byte sequentially using AVX instructions, which we replicated in Rust using SIMD instructions.
-On the FPGA side, the hash computation is parallelized and the iteration over the element’s bytes is pipelined.
-We present the numbers of using a single thread on CPU and a single compute unit on FPGA in order to keep the comparison fair.
-With the Bloom filter integrated into Timely Dataflow as a single offloaded operator, most of the additional overhead from processing sub-graphs is eliminated and the flow of data between FPGA and CPU is reduced to the minimal practical batch.
-In addition, the computational load is significantly higher compared to a simple filter.
-The $k$ hash functions consist of sequential bit shifting, addition, and XOR-ing, which are very fast in FPGA logic and deliver the offloading benefit comes from.
-Therefore in this application we concentrate on minimal practical batches and direct offloading benefits.
+In the Fleet C++ software-only baseline, the hashes were computed in parallel for each byte sequentially using AVX instructions.
+We re-implemented this in Rust as a Timely operator using Arm SIMD instructions.
+On the FPGA side, the hash computations are parallelized and iteration over the element’s bytes is pipelined.
+We measure a single CPU thread and compare it with a single offloaded operator.
 We implement the Bloom filter on the FPGA to take 128-byte elements with $k=8$ hash functions that take byte-wide inputs.
@@
 The return value is the set of 8 64-bit hashes for each element.
+In contrast to the synthetic benchmark above, this scenario has much lower progress tracking overhead (only a single operator is offloaded, compared with an entire sub-graph) and much higher computation cost (the Bloom filters themselves).
+The $k$ hash functions consist of sequential bit shifting, addition, and XOR-ing, which are fast in FPGA logic and deliver the offloading benefit even for small batches.
 Instead of directly returning the hash results, the FPGA could have computed the lookup table positions or performed the lookup itself, returning either $k$ indices or binary values.
@@
 In this experiment, we assume that the lookup table is located elsewhere and focus on efficiently calculating the hash function and transferring the results to the lookup table.
-In Figure 12, the latencies for PIO and CPU both start around 25 $\mu$s, due to the high overhead of streaming the input data.
-The real processing time of one element is 2.6 $\mu$s on the CPU and 1.7 $\mu$s when offloading with ECI.
-As overall processing time increases with the data size, we see the gap between CPU and PIO more clearly.
-After 8 KiB, same as in previous experiments, XDMA starts to perform better than other approaches as performance becomes dominated by the higher bulk throughput of PCIe DMA.
+Figure 12 shows the results.
+Latency for both PIO solutions and software (CPU) all start around 25 $\mu$s, due to the high overhead of streaming the input data.
+The per-element processing time is 2.6 $\mu$s on the CPU and 1.7 $\mu$s when offloading with ECI.
+As the batch size (and hence processing time) increases, the advantage of hardware offload increases, but so does the cost of PIO over PCIe.
+DMA latency barely increases, but remains significantly higher than that for ECI.
 Figure 12. Offloading Bloom filters to the FPGA
+As with the previous experiments, this shows the advantages of the efficient, low-latency communication between CPU and device that is possible using a cache-coherent interconnect, over and above both PIO and DMA over an interconnect like PCIe.
+We anticipate that a future implementation of CXL.mem 3.0, which similarly exploited cache coherence at the message level, would deliver similar benefits.
 6. Related work
@@
 It and the competing Gen-Z (Keeton, 2015) and CCIX (CCIX Consortium and others, 2024) protocols have meanwhile been either merged into or replaced by CXL (Li et al., 2023), which seems likely to be the standard interoperable protocol in the short- to medium-term future.
-However, commercial CXL hardware implementing the more recent revisions, which allow less restricted use cases, is still not widely available.
+However, commercial CXL 3.0 hardware implementing the more recent revisions has yet to appear.
 Other notable coherent interconnects include the RISC-V-specific TileLink (Berkeley Architecture Research, 2022; Terpstra, 2017; Cook et al., 2017) and NVIDIA’s NVLink 2.0 (NVIDIA, 2024; Wei et al., 2023a).
@@
 Recently, cache coherence between NUMA CPUs has been used to simulate communicating between software and a hypothetical cache-coherent NIC (Schuh et al., 2024).
-Other work (Tamimi et al., 2022) has already noted that the cache line is a better unit of transfer where small operations are expected, as in FastForward (Giacomoni et al., 2008) and Barrelfish (Baumann et al., 2009) protocols.
+Other work (Tamimi et al., 2022) has already noted that the cache line is a better unit of transfer for small operations, as in FastForward (Giacomoni et al., 2008) and Barrelfish (Baumann et al., 2009) protocols.
 More recently Concord (Iyer et al., 2023) communicates scheduling decisions between workers and the dispatcher via a polled cache line, converting worker threads from interrupt-driven to poll-mode “CPU drivers”.
@@
 The hXDP (Brunella et al., 2020) FPGA NIC work highlighted the high overhead of small PCIe transactions, and consequently performs small batch computations solely on the CPU.
-In kPIO+WC (Larsen and Ben Lee, 2015) the authors demonstrate the benefit of write combining for PCIe-based PIO.
-Compared to a traditional DMA NIC, their FPGA-based prototype shows better latency and throughput for small and medium-sized messages and comparable throughput for large messages.
+Larsen and Lee (Larsen and Ben Lee, 2015) demonstrate the benefit of write combining for PCIe-based PIO.
+Compared to a traditional DMA NIC, their FPGA-based prototype shows better latency and throughput for small-medium-sized messages and comparable throughput for large messages.
 The ThunderX-1 automatically performs combining for PCIe writes, and the results of Section 5.1 thus reflect this optimization.
+Current work (Liew et al., 2025) proposes a write-reordering unit in the device to simplify the use of fast PCIe posted writes and using Ensō (Sadok et al., 2023) DMA to transfer data from the device.
 Dagger (Lazarev et al., 2021) builds on CCI-P, the commercialized implementation of HARP’s coherent cache, to construct an FPGA-based NIC specialized for low latency RPC.
@@
 7. Conclusion and Future Work
-In this paper we have demonstrated that PIO on modern hardware outperforms DMA over a wide range of payload sizes in a number of different applications.
-For anything up to a disk block, VM page, or Ethernet Jumbo Frame, PIO offers (often dramatically) lower transfer times mostly thanks to its low fixed overheads.
-This advantage is consistent across RPC workloads, NIC packet transfer, and FPGA offloaded applications.
-This strongly contradicts the accepted wisdom, under which these specific payloads and applications are the canonical examples of DMA-suitable tasks.
-Moving beyond conventional interconnects (PCIe), we have shown that the advantage grows even greater for the emerging class of coherent interconnects.
-This is partly due to a more suitably-sized unit of transfer (the cache line) and partly due to the opportunity to interact with the coherence protocol in a carefully-optimized manner.
-Moving back to PIO makes sense in a landscape of plentiful, cheap CPU cores and fine-grained RPC-style workloads.
-This work is ongoing, and we intend to develop it further in several ways.
-Both the cache-coherent NIC and Timely Dataflow offload designs will continue to be optimized, to explore the limits of PIO-style device communication on emerging coherent interconnects.
-We will also further explore the design of appropriate programming interfaces and OS-level integration of coherent mechanisms for high-performance devices.
+It is important to regularly question long-held assumptions about systems, particularly in the light of new technological trade-offs and emerging workloads.
+We show that modern workloads can benefit from techniques that optimize for small, frequent, low-latency interactions between CPU cores and devices as much as the more traditional large, throughput-oriented interactions for which current PCIe interconnects and DMA are optimized.
+Moreover, the use of cache coherent interconnects between CPU cores and devices enables new techniques with this property, particularly when the device is able to participate in a standard cache coherence protocol at the level of individual cache transitions rather than naive coherence.
+We have demonstrated the potential of these techniques by showing a family of messaging protocols implemented over Enzian’s cache coherent interconnect, but more importantly this opens up a wide design space for the use of future coherent protocols like CXL.mem 3.0, which offers many possibilities for more efficient fine-grained communication within and between machines.